### v5

# Configurations, Tessellations and Tone Networks

Abstract
The Eulerian tonnetz, which is commonly represented by a network of tones, can be represented equivalently by a regular bipartite graph of degree three with twelve white vertices denoting major chords and twelve black vertices denoting minor chords. Three minor chords are associated to each major chord and three major chords are associated to each minor chord. This Levi graph can be realized geometrically as a configuration of twelve points and twelve lines in $\mathbb{R}^{2}$ with the property that three points lie on each line and three lines pass through each point. Interesting features of the tonnetz, such as the four hexatonic cycles and the three octatonic cycles, crucial for the modern understanding of nineteenth-century harmony and voice leading, can be read off directly as properties of this configuration $\{12_{3}\}$ and its Levi graph. Analogous tone networks along with their Levi graphs and configurations can be constructed for pentatonic music and twelve-tone music. These and other new tonnetze offer the promise of new methods of composition. If the constraints of the Eulerian tonnetz are relaxed to allow variations between major and minor triads at exactly two tones, the resulting bipartite graph has two components, each generating a tessellation of the plane, of a type known to Kepler, based on hexagons, squares and dodecagons. When the same combinatorial idea is applied to tetrachords of the ‘Tristan’ genus (dominant sevenths and half-diminished sevenths) the cycles of the resulting bipartite graph are sufficiently ample in girth to ensure the existence of a second configuration $\{12_{3}\}$, distinct from the Eulerian tonnetz as an incidence geometry, which can be used for a new approach to the analysis of the rich tetradic harmonies of the nineteenth century common practice.
Keywords: Music and mathematics, tonnetz, tessellations, self-dual configurations, Levi graphs,
pentatonic scale, Desargues configuration, twelve-tone system, Cremona-Richmond configuration, Wagner, Tristan genus, dominant sevenths, half-diminished sevenths, Tristan und Isolde, Götterdämmerung, Parsifal.
I Introduction
The well-known Eulerian tonnetz, which consists of an intricate network of relations between the twelve notes of the chromatic scale, was originally introduced by the mathematician Leonhard Euler (1739) and later developed in works of Gottfried Weber (1824), Ernst Naumann (1858), Arthur von Oettingen (1866) and Hugo Riemann (1902).  [^1] We treat the word ‘tonnetz’, spelled with a lower case ‘t’, and its plural ‘tonnetze’ as words of English.   Investigation of the tonnetz from a modern standpoint in the setting of equal temperament was carried forward in works of Waller (1978), Lewin (1982, 1987), Hyer (1989), Cohn (1996, 1997, 1998), Douthett & Steinbach (1998) and many others, attracting much attention. Figure 1 shows an alternative representation of the Eulerian tonnetz as a tessellation of the plane by regular hexagons in which the vertices are the conventional major and minor chords. This figure, viewed as an infinite graph, is familiar to music theorists – see, for example, Douthett & Steinbach (1998) at p. 248 or Tymoczko (2011b) at p. 413. But the tessellation in Figure 1 is more than a mere graph or diagram: it is a geometric object in the Euclidean plane.
We propose here to introduce another geometric representation of the tonnetz – specifically, we show that the tonnetz can be visualized as a configuration of twelve lines and twelve points in $\mathbb{R}^{2}$, with the property that three of the points lie on each line and three of the lines pass through each point. The construction proceeds in several steps. First, we interpret the tonnetz as a regular bipartite graph of degree three and girth six with the two types of vertices corresponding to major and minor triads, respectively, and the edges encoding adjacency within the tonnetz. This allows for easy representation and analysis of various cyclic chord progressions that arise naturally in the tonnetz. We go on to realize this Levi graph as an incidence geometry consisting of twelve points (major triads) and twelve lines (minor triads) in the Euclidean plane, in the form of a configuration $\{12_{3}\}$, leading to interesting visualizations. Finally, we identify this configuration as the self-dual $\{12_{3}\}$ of type D222 drawn and described by Daublebsky von Sterneck (1895).
The tonnetz is widely regarded as essential to the modern understanding of music based on triadic harmonies – so it is both unexpected and important to see such a connection between the tonnetz and a basic construction in Euclidean geometry. But there are challenges as well. Can the idea be generalized to other systems of music? What about pentatonic music, for example? Or the twelve-tone system? What about the richer four-note harmonies of nineteenth century common practice used in abundance by composers such as Chopin, Wagner and Tchaikovsky? The answer to these questions, as we shall see, is ‘affirmative’. Mathematicians have been studying the geometry of configurations for centuries and there is a vast literature relating to the subject. By a configuration $\{m_{r},n_{k}\}$ in $\mathbb{R}^{2}$, we mean a collection of $m$ points and $n$ lines in the Euclidean plane such that $r$ of the $n$ lines are incident with each point and $k$ of the $m$ points lie on each line. In the case of a ‘balanced’ configuration, for which the number of points is the same as the number of lines, and where the number of points lying on each line is the same as the number of lines passing through each point, we write $\{n_{k}\}$ in place of $\{n_{k},n_{k}\}$.
The triangle is a balanced configuration of type $\{3_{2}\}$ formed of three points and three lines, with two points on each line and two lines through each point. The quadrangle $\{4_{2}\}$, the pentagon $\{5_{2}\}$, the hexagon $\{6_{2}\}$, and more generally the $n$-gon $\{n_{2}\}$ can be defined similarly.  [^2] The most well known nontrivial balanced configuration in the plane is Pappus’s $\{9_{3}\}$. By ‘nontrivial’ we mean that (unlike, say, the triangle or the quadrangle) it is not a priori obvious that it should exist. The existence of this $\{9_{3}\}$ relies on Pappus’s theorem, which states that if three points $a$, $b$, $c$ lie on a line $A$ and three more distinct points $d$, $e$, $f$ lie on a line $B$ that is distinct from $A$, then if we let $x$ denote the intersection of the lines $ae$ and $bd$, let $y$ denote the intersection of the lines $af$ and $cd$, and let $z$ denote the intersection of the lines $bf$ and $ce$, then there exists a line $C$ containing $x$, $y$, $z$. That gives nine lines altogether meeting the nine points $a,b,c,d,e,f,x,y,z$ in a $\{9_{3}\}$, which is called the Pappus configuration.   These configurations arise as the subconfigurations associated with the subgraphs of the Levi graph of the tonnetz corresponding to the musical progressions derived from hexacycles, octacycles, decacycles, dodecacycles, and $2n$-cycles, respectively. There is also the notion of a ‘complete quadrangle’, which is a configuration that consists of four points in general position and the six lines joining them in pairs. The complete quadrangle is a configuration of type $\{4_{3},6_{2}\}$, since two points lie on each line and three lines pass through each point. By four points in ‘general position’ in $\mathbb{R}^{2}$ we mean that no three lie on a line and no two of the lines joining them are parallel. The complete pentagon, which consists of five points in general position and the ten lines joining them in pairs, is a configuration of type $\{5_{4},10_{2}\}$. The complete hexagon is a configuration of type $\{6_{5},15_{2}\}$. These last two we meet in connection with pentatonic music and twelve-tone music. The complete $n$-gon is a configuration of type $\{n_{n-1},\mbox{$\textstyle\frac{1}{2}$}n(n-1)_{2}\}$.  [^3] We also have the dual configurations $\{6_{2},4_{3}\}$, $\{10_{2},5_{4}\}$, $\{15_{2},6_{5}\}$ obtained by interchanging the roles of points and lines. The triangle $\{3_{2}\}$ is self-dual.
[…]
Our central thesis is that every configuration gives rise to an abstract system of music, in which there are two sets of fundamental tone clusters (analogous to the usual major and minor triads) where certain classes of transitions between chords of one type and the other are governed by the incidence relations of the configuration. We shall be mainly concerned with balanced configurations, for which the number of chords of each type is the same, and we give examples of music systems that arise in this way. In the case of the Eulerian tonnetz, one can represent the twelve major chords by a set of points in $\mathbb{R}^{2}$ and the twelve minor chords by a set of lines in $\mathbb{R}^{2}$ in such a way that they form a $\{12_{3}\}$. Thus, the tonnetz has the structure of an incidence geometry, and we are able to identify exactly which incidence geometry it turns out to be. One can then ask, following the program outlined above, whether other types of music can be represented as incidence geometries. We give several examples, the first being a geometry of type $\{10_{3}\}$, the Desargues configuration, consisting of ten points and ten lines – in our scheme representing the ten two-note clusters and the ten three-note clusters taken from a pentatonic scale. One can think of the two-note clusters and three-note clusters as being the ‘major’ and ‘minor’ chords of pentatonic music.  [^4] Our use of the terms ‘major’ and ‘minor’ here is for convenience, to mark the combinatorial similarity that arises naturally here to the corresponding structures in conventional Western music. As we show in Section VI, one can also consider situations for which single-note clusters are paired with multiple-note clusters. For instance, the twelve notes of the chromatic scale, regarded as single note clusters, together with the major chords, regarded as three-note clusters, form a $\{12_{3}\}$, which turns out to be a D222.
Our second example is based on Sylvester’s theory of duads and synthemes, which can be used to construct a geometry of type $\{15_{3}\}$ formed of fifteen lines and fifteen points, known as the Cremona-Richmond configuration. In our scheme the underlying set from which the duads (unordered pairs) and synthemes (unordered triples of mutually exclusive duads) come is a hexachord in the sense of Schönberg, that is to say, an unordered collection of six notes taken from a twelve-tone row.  [^5] One can ask whether a syntheme – an unordered triple of unordered pairs of distinct pitches drawn from a set of six – can be interpreted as a ‘chord’. Now, even in the common practice period, the notion of a chord admits a range of interpretation. We all agree that if several notes are played simultaneously on one instrument for the same period of time, then that constitutes a chord (possibly dissonant). But if the same notes are played as an arpeggio, as on a piano or harp, then in many contexts this can (and indeed should) too be regarded as a chord. The arpeggiation of chords on the violin comes as a necessity, on account of the arch of the bridge; and on the guitar, chords are strummed. Even if the notes are spread over time, and over a range of voices and instruments, and in different octaves, as long as the essential unity and identity of its elements and the gesture that draws them together can be perceived, then we can call it a chord. In the case of a syntheme, we adopt the same idea. If three pairs of duads forming an entity can be recognized as a unity, then it is a syntheme, regardless of the vagaries of its presentation. In this sense, the syntheme can be regarded as a natural generalization of the usual idea of a chord.
We presume some familiarity on the part of the reader with the Eulerian tonnetz, and indeed when we refer to ‘the tonnetz’ without further qualification, we usually mean the tonnetz of Euler, von Oettingen and Riemann in its modern form based on equal temperament. To establish some terminology and notation we develop the relevant ideas ab initio and we present a brief construction of the tonnetz that only involves elementary ideas.
We fix the chromatic scale and give the notes their usual names $C,C_{\sharp}/D_{\flat},D,D_{\sharp}/E_{\flat}$, and so on, making no distinction between enharmonic pairs. We use curly brackets (braces) to denote an unordered set of notes. Square brackets denote an ordered set or sequence of notes. Round brackets (parentheses) are used for sequences of chords. Angled brackets are used in the case of a cycle of distinct chords, in which case the final written chord repeats the initially written chord, to signal that the sequence closes. We use the term ‘duad’ for an unordered pair of distinct notes, for example $\{C,E\}$. We write $\{C,E,G\}$ to represent an unordered triad of the notes $C$, $E$, $G$, each of which might be located within any octave. In such a context, $C$ denotes an equivalence class of notes modulo octaves. Thus, we write $CM=\{C,E,G\}$ for the ‘abstract’ $C$-major chord (i.e., the chord of pitch classes).
When we write $CM=[C,E,G]$ for an ordered triad, it is usually with a specific choice of pitch in mind for the three notes, dictated by context. For the first inversion we write $CM^{(1)}=[E,G,C]$ and for the second inversion $CM^{(2)}=[G,C,E]$. Similarly, we write $Cm=[C,E_{\flat},G]$, $Cm^{(1)}=[E_{\flat},G,C]$ and $Cm^{(2)}=[G,C,E_{\flat}]$. In a situation where the inversions are important, we write $CM^{(0)}=[C,E,G]$ for the chord in its root position. There are three further permutations of a triad, $[E,C,G]$, $[G,E,C]$ and $[C,G,E]$, making six forms altogether.  [^6] The three ‘odd’ permutations are not usually given separate names, though the perceived ‘colour’ of a chord can be sensitive to a permutation of its elements. The term ‘first inversion’ is usually applied to both of the chords with $E$ in the bass position, viz. $[E,G,C]$ and $[E,C,G]$; the term ‘second inversion’ is applied to $[G,C,E]$ and $[G,E,C]$, with $G$ in the bass. Inversions often occur in the context of four-voice part writing, where the triad, in one of its six arrangements, will have one of the voices doubled.
In the case of four-note chords one has twenty-four permutations of the way in which the four notes can be distributed over the four voices. The minor sixth $Cm^{\,6}$, for example, given by $C,E_{\flat},G,A]$ in root position, admits, among its twelve odd permutations, the form $A,E_{\flat},G,C]$. The significance of this permutation is more apparent if we transpose the root chord down by a major third to give the $A_{\flat}m^{\,6}=[A_{\flat},C_{\flat},E_{\flat},F]$ in root position and its much-discussed odd permutation $F,C_{\flat},E_{\flat},A_{\flat}]$ which, in Wagner’s notation $F,B,E_{\flat},A_{\flat}]$ we recognize as the ‘Tristan’ chord in the form it first appears in the opening bars of the eponymous opera. The permutation is $1,2,3,4]\to[4,2,3,1]$. The same chord, at the same pitch levels, appears prominently, in the left hand, as an arpeggio, fortissimo, in a passage of Chopin’s $G$-minor Ballade, Opus 23, just before the più animato, in the enharmonically equivalent representation $E_{\sharp},B,D_{\sharp},G_{\sharp}]$, obtained by applying the same permutation to the $G_{\sharp}m^{\,6}$ chord.
We have segued on to the consideration of minor sixths here to stress the importance of the group of permutations of $n$ objects in chord formation.   [^7] See Callender et al (2008). We note that under the permutation $1,2,3,4]\to[4,1,2,3]$ a minor sixth becomes a half-diminished seventh. For example, the $Am$ half-diminished seventh $A^{\phi 7}=[A,C,E_{\flat},G]$ is the third cyclic inversion of $Cm^{6}$. Note also that under the permutation $1,2,3,4]\to[2,3,4,1]$ a minor seventh becomes a major sixth. For example, the first cyclic inversion of $Cm^{7}=[C,E_{\flat},G,B_{\flat}]$ is $E_{\flat}M^{6}=[E_{\flat},G,B_{\flat},C]$.  We shall have more to say on the Tristan chord later.
With these conventions at hand we turn to the construction of the Eulerian tonnetz. We observe that a major chord can be changed to a minor chord in three different ways by fixing two of the tones and altering the third. For example, in the case $CM^{(0)}=[C,E,G]$, we can lower the $C$ by a semitone to obtain $Em^{(2)}=[B,E,G]$. If we lower the $E$ by a semitone we obtain $Cm^{(0)}=[C,E_{\flat},G]$. If we raise the $G$ by a tone we obtain $Am^{(1)}=[C,E,A]$.
Similarly, a minor chord can be changed to a major chord in three different ways by fixing two of the tones and altering the third. For example, in the case of the chord $Cm^{(0)}=[C,E_{\flat},G]$, we can lower the $C$ by a tone to obtain $E_{\flat}M^{(2)}=[B_{\flat},E_{\flat},G]$ or we can raise the $E_{\flat}$ by a semitone to obtain $CM^{(0)}=[C,E,G]$ or we can raise the $G$ by a semitone to obtain $A_{\flat}M^{(1)}=[C,E_{\flat},A_{\flat}]$.
These are the only major chords that can be obtained from $Cm$ by fixing two of the tones and altering the third. It follows that associated to each unordered major chord there are three unordered minor chords, each sharing two of the pitches of the major chord; and associated to each unordered minor chord there are three unordered major chords, each sharing two of the pitches of the minor chord. We do not assume a priori that the admissible transformations between major and minor chords are of a tone or semitone in magnitude; this is an emergent feature that arises as a byproduct of our assumption that each transformation should preserve two of the three tones.
In this paper we highlight – for the first time, we believe – the unifying role of Levi graphs and configurations (both geometric and combinatorial) in the analysis of music and in the creation of new forms of music. The geometry of tessellations also plays an important part. In Section II we show that the Eulerian tonnetz can be represented by a finite bipartite graph of degree three and girth six.  [^8] Here we follow the approach of Waller (1978). This work seems to have been neglected by music theorists and we have not seen Waller’s bipartite graph discussed elsewhere in the literature of the tonnetz (but see Cohn 2012 at p. 18). There is, of course, no mention of the tonnetz in Waller’s paper itself – for at that point the tonnetz had not become an object of renewed interest in the community of music theorists, which began with the work of Lewin and his school in the 1980s – nor does Waller mention that he is working with a Levi graph and that there should be an associated configuration. This is consistent with the fact that, with a few exceptions, such as the book of Hilbert & Cohn-Vossen (1932) and the papers of Coxeter (1950, 1958), the theory of configurations, which had been active in the nineteenth century, had been going through a dormant phase, lasting from 1910 up to 1990 (Grünbaum 2009 at pp. 8-14). Derek Waller, a British mathematician who worked in category theory, algebra and graph theory, died in 1978 at the age of 37. The connection between tonnetze and configurations is our idea, as is the realization that Waller’s construction, i.e. what we now recognize as the Eulerian tonnetz, is the Levi graph of the D222 of Daublebsky von Sterneck.  The result is stated in Proposition 1 and the graph is drawn in Figure 2. Each white vertex of this bipartite graph abuts three black vertices; each black vertex abuts three white vertices. The white vertices represent major chords; the black vertices represent minor chords. We show how properties of the four mutually disjoint hexatonic cycles and the three mutually disjoint octatonic cycles can be understood in the setting of this Levi graph and we comment on the role of other cycles that arise naturally in the graph, which we illustrate in Figure 3 with several examples.
In Section III we show how the tonnetz can be regarded as a configuration. This is achieved by treating the transformations between major and minor chords as incidence relations in a combinatorial geometry. We reach the surprising conclusion that the tonnetz can be realized geometrically in the Euclidean plane $\mathbb{R}^{2}$ as a certain self-dual configuration of type $\{12_{3}\}$. Specifically, the tonnetz turns out to be the configuration $\{12_{3}\}$ of Daublebsky von Sterneck type D222. The result is stated in Proposition 2 and the configuration of major and minor chords is drawn in Figure 4.
A central argument that we maintain throughout the discussion is that the tonnetz is a mathematical object – in this case, specifically, a combinatorial configuration – that can be represented by a Levi graph. Thus, a certain abstract structure manifests itself as (i) a combinatorial configuration, (ii) a geometric configuration – and (iii) a ‘musical’ configuration, the latter being of course the tonnetz. As an application of our analysis of the cycles of the tonnetz, we take a look in Section IV at the problem of the resolution of the ‘Tristan’ chord as it arises variously in works by Wagner, Chopin and Tchaikovsky, excerpts of which we exhibit in Figures 5, 6, 7 and 8. Our preliminary analysis of compositions by these nineteenth century composers involves a reduction from tetradic harmony to triadic harmony. Such a reduction, which can be regarded as a kind of reconnaissance, a survey of the landscape before entering into tetradic territory, can all the same be surprisingly effective and numerous insights can be gained on the basis of this approach.
The idea that tetradic harmonies can be analyzed, albeit tentatively, by use of triadic tonnetz relations can be found in the works of Lewin, Cohn, Tymoczko, and numerous other authors – and to that extent indeed we but follow in their footsteps – nonetheless, the idea of using the Levi graph associated with the tonnetz configuration as the fundamental basis for such analysis is entirely new, as best as we can say, and there are three immediate benefits: (a) the discussion can often be simplified and/or made more transparent by use of the Levi graph as a very efficient ‘road map’ of the tonnetz, (b) the geometry of the chord space can be quite unambiguously visualized and surveyed, and (c) the approach can be generalized in a satisfying and unexpected way to tetradic harmonies, which we discuss in Section X.
In Section V we consider aspects of pentatonic music. Our starting point is the complete pentagon $\{5_{4},10_{2}\}$ associated with a set of five points in general position in $\mathbb{R}^{2}$, depicted in Figure 9, which leads us in Figures 10 and 11 to show how a tonnetz for pentatonic music can be modelled by use of the self-dual Desargues configuration $\{10_{3}\}$. The associated Levi graph is shown in Figure 12, in two different ways, from which the cycles of the pentatonic system can be deduced. The results are summarized in Proposition 3, which lays a foundation for new types of composition in pentatonic music. A generalization of the idea is presented in Proposition 4, which offers the scope for new compositions in any scale based on an odd number of distinct notes in the octave.
In Section VI we look more closely at the relation between the Eulerian tonnetz, represented by the Levi graph in Figure 2, and the pentatonic tonnetz, represented by the Levi graph in Figure 12. These graphs bear a striking resemblance, yet whereas construction of the Eulerian tonnetz is usually motivated by the consideration of smooth voice leading, our pentatonic tonnetz was derived via set inclusion relations. One might try to argue that if tonnetze are mainly about voice-leading, then the pentatonic construction is not really a tonnetz. We argue the opposite by demonstrating that the Eulerian tonnetz itself can be derived entirely via the consideration of set-inclusion relations, as in the pentatonic situation, with no reference to voice leading, minimal work, or parsimony. Specifically, we show that the incidence relation between the twelve notes of the chromatic scale and the twelve major chords, viewed combinatorially, is that of the D222 of Daublebsky von Sterneck. No mention of the minor chords or their voice-leading relations to major chords is involved in making this assertion. The relation between the notes of the chromatic scale and the minor chords is again that of the D222. The graphs showing the relevant inclusion relations are plotted in Figure 14. These distinct note/chord graphs can be combined with that of the Eulerian tonnetz to produce a grand tripartite tonnetz given by the tessellation shown in Figure 15. The results are summarized in Proposition 5.
In Section VII we consider twelve-tone music and show how Sylvester’s combinatorial theory of ‘duads’ and ‘synthemes’ for the permutations of six objects can be used to construct a tonnetz for music based on the specification of an unordered collection of six tones. Indeed, in the context of the system of Schönberg it has long been regarded as natural to consider such hexachords as a basis for both the composition and the analysis of music. Beginning with a hexachord, one can construct a system consisting of fifteen duads (unordered pairs of tones) and fifteen synthemes (unordered triples of non-overlapping duads). These can be represented geometrically by a configuration $\{15_{3}\}$ comprising fifteen lines and fifteen points in $\mathbb{R}^{2}$ described in a paper by Richmond (1900) and discussed in detail by Baker (1925). The Levi graph representing this configuration (Tutte 1947, Coxeter 1957) is shown in Figure 16 and a geometric representation of this $\{15_{3}\}$ is drawn in Figure 17. The results are summarized in Proposition 6. In Section VIII we return to the analysis of tonal music, with a view to tackling the problem of tetrachords. The strategy, however, is first to look at the relation between major and minor triads when exactly two tones are allowed to change in a chord transition. We show that the resulting bipartite graph has two components, each containing six major chords and six minor chords. These are shown in Figure 19. Each component has girth four: hence neither is a Levi graph – but each component can be unfolded to yield the infinite graph of an Archimedean or semi-regular tessellation – a tiling of the plane by hexagons, squares, and dodecagons, for which one component is shown in Figure 20. The resulting cycles include some interesting sonorities and in Figure 21 we present a short fanfare that we have composed based on one of the hexacycles of this tonnetz.
In Section IX we take a break from the main argument and present a short dialogue involving three friends, Salviati, Sagredo and Simplicio, who are interested in mathematics and music. They have agreed to meet for drinks at a music cafe in central London, but Simplicio is running late. Salviati and Sagredo enter into a conversation about the tonnetz and its mathematical representations. Then in Section X we look at the Tristan genus  [^9] We use the terminology of Cohn (2012), pp. 159-166. comprising the twelve dominant sevenths and the twelve minor sixths. Previous work in this area (Childs 1998, Douthart & Steinbach 1998, Tymoczko 2011b, Cohn 2012) has tended to look at irregular graphs, with the dominant sevenths and the half-diminished sevenths as vertices, and the edge sets including all possible ‘parsimonious’ transformations, or some large subset thereof, variously supplemented by further vertices representing additional, less consonant tetrachords, such as the six French sixths and the three diminished sevenths, and corresponding edges. The irregularities of these graphs, the large numbers of edges, the heterogeneity of the vertex constituents make one wonder whether one can do better.
We consider the situation when exactly two tones are allowed to change in a chord transition between two chords of the reductively opposite mode when we regard the minor sixth chord as a minor triad with an added dissonant sixth tone and we regard the major seventh chord as a consonant major triad with an added flat seventh tone. This gives a six to six map between the two sets of twelve chords. But this cannot give rise to a configuration, since the fundamental inequality (1) is violated. The inequality tells us that to obtain a bipartite graph of degree six with no cycles of degree less than six, the number of vertices must be no less than thirty-one – but we only have twenty four. To obtain a tractable subclass of transformations that leads to a geometry, we must cut down the number of allowed transformations. Our proposed solution to the problem, which is new as far as we can see, is as follows: we constrain the transformation class in such a way that it always retains the dissonant note and hence always changes two tones of the underlying major or minor chord. Note that the image of the dissonant tone need not (but might) be a dissonant tone of the transformed chord. The result is a three to three map between the dominant sevenths and the minor sixths with no tetracycles – a configuration $\{12_{3}\}$. We draw its Levi graph in Figure 22. One can tell at a glance that this Levi graph is distinct from that of the Eulerian tonnetz by counting the numbers of cycles of various lengths. A single discrepancy suffices to show that our Tristan tonnetz is different from the Eulerian one. But all the configurations $\{12_{3}\}$ are known, so if it is not the Eulerian tonnetz, which one is it? Surprisingly, it turns out that the Tristan tonnetz is the D228 of Daublebsky von Sterneck, which we show – with tetrachord labels attached – in Figure 23. This configuration is as memorable as the D222 and it has a large group of symmetries. We conclude Section X with an analysis of passages in Tristan, Götterdämmerung and Parsifal that allow for comparison of our method with arguments of Lewin (1996), Tymoczko (2011b) and Cohn (2012). For the Finale in Section XI we return to the Fidelio Cafe, where Simplicio has at last arrived.
II The Tonnetz as a Levi Graph
The tonnetz has been studied extensively in the literature of music and mathematics. It would be out of place to attempt a bibliography here, but we mention Lewin (1987), Hyer (1989), Cohn (1996, 1997, 1998, 2011, 2012), Douthett & Steinbach (1998), Childs (1998), Gollin (2008), Callender, Quinn & Tymoczko (2008), Crans, Fiore, & Satyendra (2009), Gollin & Rehding (2011), Tymoczko (2006, 2011a, 2011b, 2012), Catanzaro (2011), Yust (2018), Cubarsi (2024) and Rietsch (2024) as examples of recent work on the tonnetz.
Most investigations of the tonnetz hitherto have tended to develop the chicken-wire representation put forward by Douthett & Steinbach (1998) and the group-theoretic ideas that are implicit in this representation. Though not inconsistent with that point of view, we prefer a different approach to the tonnetz that allows more readily for the far-reaching generalizations required for treating forms of music other than that of the extended common practice period.
Our starting point is the observation that the tonnetz can be represented by a regular bipartite graph of degree three and order 24 of the form shown in Figure 2. Major chords are represented by white vertices and minor chords are represented by black vertices. Each major chord is connected to the vertices of three minor chords and each minor chord is connected to the vertices of three major chords. For example, $CM$ is connected to $Em$, CmCm, AmAm, whereas $Cm$ is connected to $E_{\flat}M$, $CM$, $A_{\flat}M$.
Various well-known features of the tonnetz can be read off from this graph and less well-known aspects of the tonnetz also become apparent. To develop our approach further, we recall a few ideas from graph theory (Harary 1969, Wilson 1972, Bondy & Murty 1976). We write $\mathbb{N}=\{1,2,3,.\,.\,.\,\}$ for the natural numbers and $\mathbb{Z}^{+}=\{0,1,2,.\,.\,.\,\}$ for the non-negative integers  [^10] There are diverse notations for these basic sets. We follow Welsh (1976) and Williams (1991).. A graph is a triple $G=(V,E,\pi)$ where $V$ is a non-empty finite set of elements called vertices, $E$ is a set of elements called edges and $\pi:E\to\mathcal{P}(V)$ is a map from $E$ to the set $\mathcal{P}(V)=\{\{u,v\}:u,v\in V\}$. We observe that $\mathcal{P}(V)$ is the set of unordered pairs of elements of $V$. Care is required when we consider elements of $\mathcal{P}(V)$ of the form $\{v,v\}$, $v\in V$. Since $\{v,v\}=\{v\}$ as sets, we have $\mathcal{P}(V)=\{\{u,v\}:u,v\in V,u\neq v\}\cup\{\{v\}:v\in V\}$. If $\pi(e)=\{v_{1},v_{2}\}$, where $v_{1},v_{2}\in V$ and $e\in E$, we say that $v_{1}$ and $v_{2}$ are joined by $e$ and that $e$ is incident with $v_{1}$ and $v_{2}$.
If $e_{1},e_{2}\in E$ are such that $\pi(e_{1})=\{v_{1},v_{2}\}$ and $\pi(e_{2})=\{v_{1},v_{2}\}$, where $v_{1},v_{2}\in V$, then we say that $e_{1}$ and $e_{2}$ are multiple edges. A graph $G=(V,E,\pi)$ has no multiple edges if and only if $\pi$ is injective. If $e\in E$ is an edge such that $\pi(e)=\{v,v\}$ for some $v\in V$, then $e$ is a loop at $v$. If $e_{1},e_{2}\in E$, $e_{1}\neq e_{2}$, are such that $\pi(e_{1})=\{v,v\}$ and $\pi(e_{2})=\{v,v\}$ for some $v\in V$, then we say that $e_{1}$ and $e_{2}$ are multiple loops at $v$.
The number of edges joining a vertex is the degree of that vertex. The degree of a graph is the maximum of the degrees of its vertices. If all the vertices have the same degree, the graph is regular. If the edges $e_{1}$ and $e_{2}$ are incident with a common vertex, we say that $e_{1}$ and $e_{2}$ are adjacent. If a vertex is incident with no edges it is isolated. If $G$ has no loops or multiple edges it is said to be simple. A sequence of $k\in\mathbb{N}$ distinct consecutively adjacent edges is a path of length $k$. A graph is connected if any two vertices are joined by some path. A subgraph of $G=(V,E,\pi)$ is a graph $G^{\prime}=(V^{\prime},E^{\prime},\pi^{\prime})$ such that $V^{\prime}\subset V$, $E^{\prime}\subset E$, $E^{\prime}=\pi^{-1}(\mathcal{P}(V^{\prime}))$, and $\pi^{\prime}$ is the restriction of $\pi$ to $E^{\prime}$. If $G$ is not connected then it splits into two or more connected subgraphs called components of $G$.
A closed sequence of $k$ distinct consecutively adjacent edges that begins at some vertex $v$ and comes back to $v$ in such a way that all $k$ of the vertices are distinct is called a $k$-cycle. A loop is a $1$-cycle; a pair of multiple edges form a $2$-cycle. The least value of $k$ for which a graph $G$ admits a $k$-cycle is called the girth of $G$. If $V$ has $N$ elements then $N$ is the order of the graph. If the vertices of a regular graph are of degree $d$, then $dN=2M$, where $M$ is the number of edges. By a forest we mean a graph with no cycles. By a tree we mean a connected forest. Finally, we say that a graph is bipartite if there exist sets $V_{0}(G)$ and $V_{1}(G)$ satisfying $V(G)=V_{0}(G)\cup V_{1}(G)$ and $V_{0}(G)\cap V_{1}(G)=\varnothing$ such that every edge of $G$ joins a vertex of $V_{0}$ to a vertex of $V_{1}$. One can show that a graph is bipartite if and only if every cycle has an even number of edges.
We also require the combinatorial idea of an incidence structure. By an incidence structure we mean a disjoint pair of finite sets $\mathbb{P}$ and $\mathbb{L}$, whose elements we call points and lines, satisfying certain conditions. We assume that $\mathbb{P}$ and $\mathbb{L}$ are equipped with an incidence relation $R:\mathbb{P}\times\mathbb{L}\to\{0,1\}$ such that for any $p\in\mathbb{P}$ and any $L\in\mathbb{L}$ either $R(p,L)=1$, when we say that $p$ and $L$ are incident, or $R(p,L)=0$, when they are not incident. To aid intuition, we use geometrical terms such as ‘$p$ lies on $L$’ or ‘the lines $L$ and $M$ intersect at the point $p$’ to express incidence relations. The use of such language is purely for convenience; since the terms ‘line’ and ‘point’ have no a priori geometrical meaning. We assume that two distinct lines intersect at most once and that two distinct points lie on at most one line in common; thus, if $\in\mathbb{L}, LML\neq M$, then $\exists$ at most one $p\in\mathbb{P}$ such that $R(p,L)=1$ and $R(p,M)=1$; if $\in\mathbb{P}, pqp\neq q$, then $\exists$ at most one $L\in\mathbb{L}$ such that $R(p,L)=1$ and $R(q,L)=1$. We call a triple $\{\mathbb{P},\mathbb{L},R\}$ satisfying these conditions an incidence structure.
Figure 2: The Eulerian tonnetz as a Levi graph. The tonnetz can be modelled as a regular bipartite graph of degree three with twelve white vertices corresponding to major chords and twelve black vertices corresponding to minor chords. Each major chord is linked to three minor chords and each minor chord is linked to three major chords. The four $3p$-hexacycles are blunted triangles such as $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$. There are twelve $2p$-hexacycles (bow ties), one beginning at each major chord, an example being $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$. Each major chord is contained in three $2p$-hexacycles; similarly for minor chords.
By a combinatorial configuration of type $\{m_{r},n_{k}\}$ for $m,n,r,k\in\mathbb{N}$ we mean an incidence structure with $m$ points and $n$ lines such that $k$ points lie on each line and $r$ lines go through each point. We use the term ‘type’ here since in some cases several different configurations of the same type may exist. By a flag we mean a pair $\{p,L\}$ where $p\in\mathbb{P}$ and $L\in\mathbb{L}$.
For any configuration $\{m_{r},n_{k}\}$ we have the fundamental equality $mr=nk$, since each side of this equation gives a way of counting the total number of distinct flags in the configuration. The fundamental inequality
$m\geq r(k-1)+1$ (1)
follows from the fact that the total number of points $m$ must be no less than the total number of points lying on any given set of flags sharing a point in common.
If the number of lines equals the number of points in a configuration then by the fundamental equality the number of points on each line must equal the number of lines through each point. Then we say that the configuration is ‘balanced’, in the terminology of Grünbaum (2009). Since balanced configurations are frequently studied, the abbreviation $\{n_{k}\}$ is used to denote a configuration of type $\{n_{k},n_{k}\}$.
There is a special type of graph associated with any combinatorial configuration called its ‘Levi graph’ (Levi 1929, 1942, Coxeter 1950). The Levi graph $G(C)$ of a combinatorial configuration $C$ of type $\{m_{r},n_{k}\}$ is the bipartite graph that has $m$ white vertices, corresponding to points of $C$, and $n$ black vertices, corresponding to lines of $C$. A pair of vertices of $G$ determine an edge of the graph if and only if one represents a point of $C$ and the other represents a line of $C$ that is incident with that point. It follows that (i) each edge joins a white vertex and a black vertex, (ii) $r$ edges meet each white vertex, and (iii) $k$ edges meet each black vertex. Clearly, for such a graph the total number of edges is given by $mr=nk$.
The significance of Levi graphs is that their correspondence with combinatorial configurations is one to one (Grünbaum 2009, p. 28). To wit: a graph $G$ is the Levi graph of a combinatorial configuration $\{m_{r},n_{k}\}$ if and only (i) $G$ is bipartite with partition $V(G)=P(G)\cup L(G)$, where $P$ has cardinality $m$ and $L$ has cardinality $n$; (ii) every vertex in $P$ is of degree $r$ and every vertex in $L$ is of degree $k$; (iii) $G$ is simple; (iv) $G$ has girth at least six. With these definitions and facts at hand, we are able to assert the following:
Proposition 1.
The Eulerian tonnetz can be represented by the Levi graph of a configuration of type $\{12_{3}\}$ in which the $12$ major chords are represented by white vertices and the $12$ minor chords are represented by black vertices. The $36$ edges are arranged in such a way that each white vertex is connected to $3$ black vertices and each black vertex is connected to $3$ white vertices – that is to say, the graph is regular of degree three.
Inspection of Figure 2 shows the minimal cycles are hexacycles. Hexacycles can be classified by the number of parallel transformations that they contain, that is, transformations of the form $CM$ to $Cm$ or $Cm$ to $CM$. Figure 2 shows that there are exactly four hexacycles admitting three such parallel transformations, each taking the form of a sort of blunted triangle. These four distinctive hexacycles partition the tonnetz and are given as follows:
$\hskip 8.5359pt\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle,\hskip 14.22636pt\langle FM,Fm,D_{\flat}M,C_{\sharp}m,AM,Am,FM\rangle,\hskip 14.22636pt$
$\langle B_{\flat}M,B_{\flat}m,G_{\flat}M,F_{\sharp}m,DM,Dm,B_{\flat}M\rangle,\hskip 14.22636pt\langle E_{\flat}M,E_{\flat}m,BM,Bm,GM,Gm,E_{\flat}M\rangle.$ (2)
Recall that when we write out a sequence of chords forming a cycle, we use angled brackets and we repeat the initial chord at the end of the sequence to show how it closes up. One recognizes these cycles as the four mutually disjoint hexatonic cycles arising in neo-Riemannian theories (Cohn 1996, 1997, 1998, 2012, Douthett & Steinbach 1998).  [^11] See Cohn (2012) at p. 211 for a definition of ‘hexatonic cycle’. Each such hexacycle is constructed from exactly six distinct tones and contains three distinct parallel transformations represented by moving across the Levi graph.
We call these cycles thrice-parallel, which we abbreviate as $3p$. One of the advantages of the use of a Levi graph for the tonnetz is that it allows one to see at a glance the symmetries of the four $3p$-hexacycles and how they partition the tonnetz.
Figure 3: We illustrate a $3p$-hexacycle, a $2p$-hexacycle, a $4p$-octacycle and a $2p$-decacycle. The $3p$-hexacycle $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$ is one of the four basic hexatonic sequences discussed by Cohn and numerous other authors. There are twelve overlapping ‘straight bow tie’ $2p$-hexacycles. We see $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$ as an example. Similarly, we have the ‘four-cornered hat’ $4p$-octacycle $\langle CM,Am,AM,F_{\sharp}m,G_{\flat}M,E_{\flat}m,E_{\flat}M,Cm,CM\rangle$ and the ‘floppy bow tie’ $2p$-decacycle $\langle CM,Am,FM,Dm,B_{\flat}M,B_{\flat}m,D_{\flat}M,Fm,A_{\flat}M,Cm,CM\rangle$.
The edges of the tonnetz fall into three sets of twelve. In the first set we have edges of the form $\{CM,Cm\}$, which we already mentioned, that transform a major (resp., minor) chord into its parallel minor (resp., major). In the second set we have edges of the form $\{CM,Am\}$ that transform a major (resp., minor) chord into its relative minor (resp., major). In the third set we have edges of the form $\{CM,Em\}$ that transform a major (resp., minor) chord into a corresponding minor (resp., major) under a leading tone exchange. Three edges meet at each vertex, one of each type.
One can ask if there are any other hexacycles in the Levi graph of the tonnetz apart from the four hexatonic hexacycles and the answer is affirmative. There are twelve further hexacycles, one beginning at each major chord, of which a typical example is
$\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle.$ (3)
This cycle includes only two parallel transformations – viz., from $CM$ to $Cm$ and from $FM$ to $Fm$; hence, the twelve $2p$-hexacycles are distinct from the four $3p$-hexacycles. The $2p$-hexacycles are what Cohn (2012), pp. 113-121, calls ‘pitch retention loops’; at p. 43 of Cohn (1997), they are called ‘LPR loops’.  [^12] Cohn’s $LPR$-loops are not ‘loops’ in the terminology of graph theory – they are ‘cycles’.
An $L$-transformation sends a major chord one notch counterclockwise on the Levi graph to the previous minor chord, and it sends a minor chord one step clockwise to the next major chord. An $R$-transformation sends a major chord notch clockwise on the Levi graph to the next minor chord, and it sends a minor chord one step counterclockwise to the previous major chord. A $P$-transformation sends any chord across the Levi graph to its mode reverse.
The general pattern of a $2p$-hexacycle is as follows. Beginning at the chord $CM$ the cycle moves two steps clockwise along the graph to $FM$; then it crosses the graph with a $P$-transformation to $Fm$; then it moves two steps counter-clockwise to $Cm$, and finally back to $CM$ with another $P$-transformation. From left to right the string of operations applied to the initial chord $CM$ in this example going clockwise thus takes the form $RLPRLP$, or else $PLRPLR$ going counterclockwise.
The other remaining eleven $2p$-hexacycles can be obtained by moving the initial point around in a circle of fourths, that is to say, from $CM$ to $FM$ to $B_{\flat}M$ and so on. Each tone is contained in three distinct $2p$-hexacycles. The $3p$-hexacycle (‘blunted triangle’) beginning at $CM$ is generated by the string $PLPLPL$ clockwise and the string $LPLPLP$ counterclockwise. This is the unique hexatonic cycle containing that chord. The $1p$-octacycle (‘octabeanie’) beginning at $CM$ is generated by the string $RLRLRLRP$. And the $4p$-octacycle (‘four-cornered hat’) beginning at $CM$ is generated by $RPRPRPRP$ going one way, or else $PRPRPRPR$ going the other.
III The Tonnetz as a Configuration
‘. . . interesting configurations are represented by interesting graphs.’
—H. S. M. Coxeter, Twelve Geometrical Essays
If we accept the view held by many that Euler’s tonnetz is not a mere musico-mathematical curiosity but is fundamental to the structure of music (specifically, to that of the common practice period), then it makes sense to look at the mathematics of the tonnetz from every angle – group-theoretic, number-theoretic, set-theoretic, combinatoric and geometric – for insights that might be gained.
One key concept, lying on a nexus of the various branches of mathematics mentioned above, is the idea of configuration that we introduced in Section I. By a configuration, in the broadest sense, by which we mean combinatorial, we mean a set of $m$ points and $n$ lines such that every point lies on exactly $r$ lines and every line contains exactly $k$ of these points. Some variants of the idea arise when we consider the interpretations of the words ‘point’ and ‘line’ and when we consider the spaces in which these elements are taken.
Configurations of the form $\{n_{3}\}$ have attracted much attention and have been investigated at length. The inequality (1) shows that no $\{n_{3}\}$ configurations exist for $n\leq 6$. We draw on this fact a little later. It can be interesting to take stock of some of the configurations that exist for higher values of $n$. There is a unique configuration $\{7_{3}\}$, modelled by the seven points and seven lines of the finite projective plane based on the ‘binary’ field $\mathbb{Z}_{2}$. This is the Fano plane. There is a unique combinatorial configuration of type $\{8_{3}\}$, which can be realized explicitly by a system of eight points and eight complex projective lines in the complex projective plane $\mathbb{CP}^{2}$. No real coordinatization of $\{8_{3}\}$ on the Euclidean plane exists. There are three distinct combinatorial configurations $\{9_{3}\}$, all of which can be modelled as systems of points and lines in $\mathbb{R}^{2}$. One of these is the Pappus configuration mentioned earlier. All three of the configurations $\{9_{3}\}$ can be modelled over the rational numbers. There are ten distinct combinatorial configurations $\{10_{3}\}$, nine of which can be modelled as geometric configurations in $\mathbb{R}^{2}$ and are rationally realizable. Among these is the famous Desargues configuration $\{10_{3}\}$, about which we shall have more to say later. But one of the combinatorial configurations $\{10_{3}\}$ does not admit a representation as a geometric configuration over any field.
The 31 distinct configurations of type $\{11_{3}\}$ were identified by Daublebsky von Sterneck (1894), all of which can be realized in $\mathbb{R}^{2}$.  [^13] In his paper on the configurations $\{11_{3}\}$, Daublebsky von Sterneck (1894) says the following, which we can presume applies equally to his later work on configurations $\{12_{3}\}$: ‘Es handelt sich hier nur darum, alle möglichen Schemata, welche obige Eigenschaften besitzen, herzustellen und ist die Aufgabe insoferne ein Problem der Combinationslehre. Die im Folgenden verwendeten geometrischen Betrachtungen dienen bloss dazu, die combinatorischen Überlegungen zu unterstützen und zu erleichtern.’ In English: ‘Here the only goal is to create all possible systems which possess the above $\{11_{3}\}\,]$ properties and in this respect the task is a problem of combinatorics. The geometric considerations used in the following serve merely to support and clarify the combinatorial considerations.’ The remark is interesting in the present context, since one can take the view that the tonnetz is primarily a combinatoric object, well-represented as such by its Levi graph, and that its geometric representation as a configuration in $\mathbb{R}^{2}$ mainly serves to support and clarify the combinatorial considerations. This is the minimal position that one can take with reference to the role of combinatorial configurations in music. Nonetheless, combinatorics and geometry have come a long way since the days of Daublebsky von Sterneck and from a modern perspective we take a more unified approach to the relations between these disciplines – and hence the view that music is somehow deeply entangled with this unity. The test as to whether this is really the case is to see whether entertaining such a view is fruitful in the new ideas that it generates in the context of mathematical music theory. The situation with $\{12_{3}\}$ is not completely straightforward. Daublebsky von Sterneck (1895, 1903) established the existence of 228 distinct configurations of this type. This number was accepted as definitive as recently as the work of Sturmfels & White (1990). Then Gropp (1991, 1993, 1997) showed that one case was missing from Daublebsky von Sterneck’s list: hence, in fact, the number of $\{12_{3}\}$ combinatorial configurations is 229. Sturmfels & White (1990) were investigating the conjecture of Grünbaum that every $\{n_{3}\}$ that can be realized over the reals can be realized over the rationals. They established by use of diophantine computations that all the $\{11_{3}\}$ and $\{12_{3}\}$ cases considered by Daublebsky von Sterneck were realizable over $\mathbb{Q}$. Thus it follows from a long line of work by geometers and combinatorialists spanning more than a century that all 229 of the combinatorial $\{12_{3}\}$ configurations admit geometric realizations in $\mathbb{R}^{2}$ and indeed in the rational plane.  [^14] See Betten, Brinkmann & Pisanskic (2000), Gropp (2004), Grünbaum (2009) and Alazemi & Betten (2014) for details of the history of the subject and for tables and constructions for $\{12_{3}\}$ and $\{n_{3}\}$ for higher values of $n$.
Now we are in a position to say how the tonnetz fits into this picture. We know by inspection of Figure 2 that the tonnetz can be represented as the Levi graph of a combinatorial configuration of type $\{12_{3}\}$. But we also know that each such combinatorial configuration $\{12_{3}\}$ can be realized as a geometric configuration $\{12_{3}\}$ in $\mathbb{R}^{2}$. That implies, as a consequence, that the tonnetz is such a configuration – that it can be drawn as a geometric configuration of points and lines in the Euclidean plane, the major chords being represented by points and the minor chords by lines, so that the points lie on lines in threes and the lines go through points in threes.
But which configuration? Once one gets this far, the problem that remains is to determine which one of 229 known configurations of type $\{12_{3}\}$ is the Eulerian tonnetz. We have been able to identify it as the configuration D222. This remarkable configuration of points and lines, labelled with major and minor chords, which we call the tonnetz configuration, is depicted in Figure 4, in line with the original drawing of Daublebsky von Sterneck (1895). Thus, we have obtained a truly interesting and unexpected link between music and geometry! Summing up, we have the following:
Proposition 2.
The Eulerian tonnetz can be represented in $\mathbb{R}^{2}$ as a self-dual configuration $\{12_{3}\}$ of Daublebsky von Sterneck type D222. The twelve major chords are represented by points and the twelve minor chords are represented by lines. Three of the points lie on each line and three of the lines pass through each point. The thirty-six incidence relations between the twelve points and the twelve lines determine the edges of the associated Levi graph.
Figure 4: The Eulerian tonnetz admits a realization as a geometric configuration on the Euclidean plane with a $\{12_{3}\}$ incidence scheme of Daublebsky von Sterneck type D222. Major chords are represented by points and minor chords by lines. Through each point representing a major chord one finds three lines representing the minor chords associated with it; and on each such line representing a minor chord one finds three points representing the three associated major chords. The four $3p$-hexacycles such as $\langle CM,Cm,A_{\flat}M,G_{\sharp}m,EM,Em,CM\rangle$ form a closed sequence of four triangles, which can be easily identified, with the property that the vertices of each such triangle are inscribed in the lines of the next. The remaining twelve triangles of the configuration, which correspond to 2pp-hexacycles, can also be readily seen and counted.
IV Cyclic Structure as a Basis for Musical Analysis
Clearly, any geometric configuration gives rise to a combinatorial configuration. It follows that any geometric configuration gives rise to a Levi graph, the white vertices corresponding to points and the black vertices to lines. Thus a triangle in $\mathbb{R}^{2}$ determines a hexagonal Levi graph. In particular, the hexacycles of the tonnetz are represented by triangles of the tonnetz configuration in $\mathbb{R}^{2}$. The 3pp-hexacycle containing the chord $CM$, for example, can be represented by a triangle in which the three major chords $CM$, EMEM, $A_{\flat}M$ correspond to points in $\mathbb{RP}^{2}$ and the three minor chords $Em$, $C_{\sharp}m$, $Cm$ correspond to the lines that join them. But the 2pp-hexacycles also correspond to triangles in the tonnetz configuration. For example, the 2pp-hexacycle $\langle CM,Am,FM,Fm,A_{\flat}M,Cm,CM\rangle$ is the triangle with vertices $CM,FM,A_{\flat}M$. It is then a simple visual exercise to count the total number of triangles in Figure 4, allowing us to determine altogether exactly sixteen different hexacycles in the tonnetz. The same exercise can be carried out in the Levi graph, since its structure is isomorphic to that of the configuration, but the counting is not so easy; and if one attempts this exercise in the setting of the traditional chicken-wire representation of the tonnetz, such as that shown in Figure 1, one can easily get lost, running about like a headless chicken. The point is that each of these representations of the tonnetz – tessellation, Levi graph, configuration – has merit in revealing aspects of the structure of the tonnetz. Here we have a good example of how the geometric structure of the tonnetz as a configuration in $\mathbb{R}^{2}$ makes is easy to identify certain key elements of musical structure – namely, the hexacycles – which are but the humble triangles of the configuration.
Similarly, one sees that the octacycles of the tonnetz are represented by quadrilaterals of the tonnetz configuration. The 4pp-octacycle $\langle Cm$, $E_{\flat}M$, $E_{\flat}m$, $G_{\flat}M$, $F_{\sharp}m$ AMAM, AmAm, $CM\rangle$ arising here as a blade-like quadrilateral with vertices $E_{\flat}M,G_{\flat}M,AM,CM$ may look familiar. It is identical to the chord sequence of the ‘four-cornered hat’ starting at $Cm$ in Figure 2. But it is also identical to the chord sequence arising in Schubert’s Overture to Zauberharfe (see Cohn 2012, pp. 85-89). The symmetric beauty of Schubert’s composition seems to be associated with the fact that it is based on a 4pp-octacycle. In particular, the three 4pp-octacycles, which correspond to a system of three mutually exclusive blade-like ‘box-cutter’ quadrilaterals in the configuration, partition the tonnetz and Schubert’s trajectory in the Overture to Zauberharfe is confined to one of these sequences.  [^15] There is no ‘musical information’ in a combinatorial configuration that is not already implicit in the associated Levi graph. This is because they are isomorphic. Nonetheless, the visualization of geometric configurations makes them ideal for the representation of certain musical ideas. A good example of this is the surprising visualization of the octacyclic progression of dominant sevenths and half-diminished sevenths in the Tristan tonnetz presented in Section X in our discussion of the final scene of Götterdämmerung.
Note that the complete quadrangle and the complete quadrilateral are ‘dual’ to each other, in the sense that the roles of the lines and the points are interchanged. The triangle, on the other hand, is ‘self-dual’ in the sense that if we interchange the roles of lines and points then we get another triangle. In the case of a self-dual configuration the number of points must be equal to the number of lines and the number of lines through each point must be the same as the number of points on each line.
It should be evident that any self-dual configuration is balanced. The triangle is thus a $\{3_{2}\}$ and for each $m\geq 3$ we can construct a self-dual configuration $\{m_{2}\}$ in the plane. On the other hand, a balanced configuration is not necessarily self-dual. This is not so obvious, since the combinatorial configurations associated with low order symmetric Levi graphs all turn out to be self-dual. In particular, among the balanced combinatorial configurations, the Fano $\{7_{3}\}$, the Möbius $\{8_{3}\}$, the Pappus $\{9_{3}\}$, the Desargues $\{10_{3}\}$, along with the other two $\{9_{3}\}$ configurations and the nine other $\{10_{3}\}$ configurations are all self-dual. But of the 31 configurations of type $\{11_{3}\}$, only 25 are self-dual, and of the 229 configurations of type $\{12_{3}\}$, only 95 are self-dual (Betten, Brinkmann & Pisanskic 2000).
As we have remarked, all the $229$ combinatorial configurations of type $\{12_{3}\}$ can be represented in the real plane and indeed the rational plane. This applies in particular to the D222 (Daublebsky von Sterneck 1895, 1903, Sturmfels & White 1990, Betten, Brinkmann & Pisanskic 2000, Alazemi & Betten 2014).
A cycle of the tonnetz represents a sequence of chords progressing via parsimonious voice leading in such a way that the sequence eventually closes upon itself. We have observed that a $2p$-hexacycle contains precisely two major/minor pairs. In such a pair, the two chords differ in the root position by a semitone in the second note. The remaining two chords in a $2p$-hexacycle differ by a semitone in the first and third notes of the chord. Thus, in the case of (3) we see that $Am$ and $A_{\flat}M$ differ by semitones in the first and third notes. One can check that each triad, major and minor, belongs to three $2p$-hexacycles. For example, $CM$ belongs to the $2p$-hexacycles beginning at $CM$, EMEM and $GM$; whereas $Cm$ belongs to the $2p$-hexacycles beginning at $CM$, $E_{\flat}M$ and $GM$.
That the twelve $2p$-hexacycles overlap in this way adds a depth of complexity to the scheme of triadic relations that may at first glance seem formidable but on reflection adds interest in a way that contrasts remarkably with the mutual disjointedness of the four $3p$-hexacycles, hence allowing for architectural modalities. As an example, we mention the possibility of ‘surgical insertion’ of one cycle within another. Suppose we consider the 2pp-hexacycle (3) alongside the transposed cycle lying a fourth above,
$\langle FM,Dm,B_{\flat}M,B_{\flat}m,D_{\flat}M,Fm,FM\rangle.$ (4)
Since the short sequence $[FM,Fm]$ appears as a subsequence of the cycle (3), we can ‘snip’ the cycles (3) and (4) at $[FM,Fm]$, and splice them together to create a 2pp-decacycle,
$\langle CM,Am,FM,Dm,B_{\flat}M,B_{\flat}m,D_{\flat}M,Fm,A_{\flat}M,Cm,CM\rangle.$ (5)
The result is not displeasing and is immediately suggestive of a new manner of composition whereby cycles can be combined to produce new cycles. For example, the procedure can be iterated, allowing one to splice in the 2pp-hexacycle
$\langle B_{\flat}M,Gm,E_{\flat}M,E_{\flat}m,G_{\flat}M,B_{\flat}m,B_{\flat}M\rangle$ (6)
by a similar snip-and-insert technique, leading to the 2pp-tetrakaidecacycle
$\langle CM,Am,FM,Dm,B_{\flat}M,Gm,E_{\flat}M,E_{\flat}m,G_{\flat}M,B_{\flat}m,D_{\flat}M,Fm,A_{\flat}M,Cm,CM\rangle.$ (7)
The surgical procedures that we have just discussed can be given a more systematic treatment by the use of the methods of graph theory, which we discuss below.
We observe also that beginning at the chord $CM$ in Figure 2, the progression moving clockwise around the perimeter of the Levi graph involves a sequence of transformations that we call ‘elevations’. More precisely, from the initial chord $CM^{(0)}=[C,E,G]$ we move to $Am^{(2)}=[C,E,A]$ by raising the $G$ to an $A$. Then we convert $Am^{(2)}$ to $Am^{(0)}$ by an action of the permutation group and perform another elevation, taking us from $[A,C,E]$ to $[A,C,F]$ by raising the note $E$ to an $F$. Then, converting the $FM^{(2)}$ to $FM^{(0)}$ by use of the permutation group and performing another elevation by raising the $C$ to a $D$, we get to $Dm^{(2)}$. In this way, we march from vertex to vertex through all twenty-four major and minor chords, giving us a Hamiltonian cycle (Harary 1969, pp. 65-70), i.e., a path that includes every vertex of the graph without repetitions. As Tymoczko (2012) remarks, a Hamiltonian path touches on all the vertices in a graph without passing through any of them twice, and is in that sense a generalization of the idea of a twelve-tone row.  [^16] See Tymoczko (2011b, 2020) on geometrical models of voice leading.
Examples of the $3p$-hexacycles and $2p$-hexacycles associated with the tonnetz can be found in Figure 3. Similarly, one can construct $4p$-octacycles and $2p$-decacycles, examples of which are also shown in Figure 3. We consider these cycles, alongside other such examples, to be architectural structures that can be used by composers in their work.
As usual with matters of artistic creativity, there are no fixed recipes for the modes in which the compositional elements should be employed; rather, merely suggestions of possibility. Nonetheless, it makes sense to let mathematics lead the way. On the matter of combining cycles together to produce new cycles, some clarity can be added if we recall the following well-known result from graph theory. Let $C_{1}$ and $C_{2}$ be cycles of a graph $G$ and suppose that the two cycles share an edge $e$ in common. Then there exists a cycle $C_{3}\subseteq\{C_{1}\cup C_{2}\}\backslash e$. That was indeed the setup, for example, with the hexacycles (3) and (6), which shared the edge $[FM,Fm]$, hence leading to a decacycle. In more detail, taking the union of the sets of the edges in the two hexacycles and removing the common edge, we were left with a set whose elements form a new cycle, a decacycle. Such operations can be pursued at greater length.  [^17] See Cohn (2012) for discussions of the cycles of the tonnetz. The idea of a cycle can be generalized by introducing the concept of a ‘cycle vector’ (Harary 1969, chapter 4). By a cycle vector in a graph $G$ we mean a collection of cycles of $G$ that are mutually edge-disjoint – that is to say, their edges do not overlap. A cycle vector is a formal sum of edge-disjoint cycles. The collection of all cycle vectors on a connected graph $G$ forms a vector space $\mathbb{V}$ over the binary field $\mathbb{F}(2)=\{0,1\}$ with the usual multiplication laws and the addition laws 0+0=0,  0+1=1,  1+1=00+0=0,\,\,0+1=1,\,\,1+1=0.
The $4p$-octacycles are of interest in their own right as objects of musical analysis on account of their relationship to octatonic scales. An octatonic scale is defined to be a scale with eight tones in it, as opposed to the seven tones of the conventional major and minor scales, with the property that adjacent notes within the scale differ either by a semitone or a tone, alternating the gap at each step. One construction of such scales is as follows. There are three diminished seventh chords, each composed of four notes. If we form the union of two of these sets of four notes, this can be done in three ways, leading to three distinct sets of eight notes. Once these notes are ordered by pitch we obtain three distinct octatonic scales. The three diminished seventh chords, $X=[C,E_{\flat},G_{\flat},A]$, $Y=[D_{\flat},E,G,B_{\flat}]$, and $Z=[D,F,A_{\flat},B]$ are disjoint as sets. Taking three distinct unions of pairs we get the three octatonic scales, $O_{12}=X\cup Y=[C,D_{\flat},E_{\flat},E,G_{\flat},G,A,B_{\flat},C]$, $O_{23}=Y\cup Z=[D_{\flat},D,E,F,G,A_{\flat},B_{\flat},B,D_{\flat}]$ and $O_{31}=Z\cup X=[D,E_{\flat},F,G_{\flat},A_{\flat},A,B,C,D]$. Any octatonic scale, starting at any chosen note, belongs to one of these three types, possibly after a cyclic permutation of its elements. Octatonic scales have been used to good effect by a number of well-known composers, among them Rimsky-Korsakov, Scriabin, Ravel, Stravinsky and Messiaen – and the use of these scales can be found in the works of earlier musicians as well (Messiaen 1956, Dallin 1964).
The relation of the octatonic scales with the tonnetz is as follows (Douthett & Steinbach 1998). Four major chords and four minor chords can be constructed from the constituents of an octatonic scale. For example, from $O_{12}$ one constructs $CM$, AmAm, AMAM, $F_{\sharp}m$, $G_{\flat}M$, $E_{\flat}m$, $E_{\flat}M$, and $Cm$. But these are the chords of the 4pp-octacycle starting at $CM$ in the tonnetz (see lower left quadrant of Figure 3). The other two 4pp-octacycles of the tonnetz can be constructed similarly from the other two octatonic scales. The 4pp-octacycles are tied to the three diminished seventh chords, in the same spirit that the four $3p$-hexacycles are tied to the four augmented triads. By taking the six possible unions of pairs of the four augmented triads we obtain six distinct hexachords. From four of these, one constructs the four $3p$-hexatonic sequences. The remaining two hexachords contain the standard whole-tone scales, each of which contains six notes. These are $W_{1}=[C,D,E,G_{\flat},A_{\flat},B_{\flat},C]$ and $W_{2}=[D_{\flat},E_{\flat},F,G,A,B,D_{\flat}]$. These arrays of six equally spaced whole notes have been exploited by Bartok, Berg, Borodin, Debussy, Janacek, Liszt, Puccini, Rimsky-Korsakov and many other composers.
Numerous examples of the use of melodic and harmonic structures based on cycles of the tonnetz in music of the common practice period have been noted (Cohn 2012 and references cited therein) in works of composers such as Mozart, Beethoven, Schubert, Schumann, Chopin, Wagner, Liszt, Bruckner, Brahms, Mahler and Strauss. Such use is not obvious to the classically trained ear since the progressions associated with tonnetz cycles do not form part of the training in basic harmony taught in everyday music theory. Nor would the composers listed have been conscious of their own use of the structural determinants implicit in the cycles of the tonnetz. If the unconscious mind will tap the resources of the mathematical structures embedded in the geometry of the tonnetz then surely it will seek to avail itself of the full richness of these resources. One advantage of the use of the twelve 2pp-hexacycles, the four 3pp-hexacycles, the twelve $1p$-octacycles, the three $4p$-octacycles, and so on, all together, in their manifestation as elements of a single cycle space, is that collectively the cycles of the tonnetz incorporate a diversity of modality sufficient to enable them to act as a basis both for the composition of music and for its analysis; thus, in our view the general scheme of pan-triadic musicology should involve all the 5409 cycles of the tonnetz. See the Appendix for an enumeration of the different types of cycles of the tonnetz.
As an example of the analysis of cyclic structure, let us return to the problem of understanding the resolution of the Tristan chord. Wagner offers a well known ‘tentative’ resolution of his permuted $G_{\sharp}m^{\,6}$ to $E^{\,7}$, as shown in Figure 5. Chopin, on the other hand, resolves the same chord to $B_{\flat}^{\,7}$, as we see in Figure 6. Chopin’s resolution is definitive: it brings that lengthy first section of the Ballade to a decisive close, and paves the way forward at the più animato with a change of tempo and the introduction of new material and a new atmosphere. Wagner’s resolution is inconclusive, in line with the ever shifting sensibilities of the protagonists.
Figure 5: Wagner, Tristan und Isolde, Act 1, opening bars. The Tristan chord $F,C_{\flat},E_{\flat},A_{\flat}]$, which in Wagner’s notation takes the form $F,B,E_{\flat},A_{\flat}]$, is one of the twelve odd permutations of the minor sixth $A_{\flat}m^{6}=[A_{\flat},C_{\flat},E_{\flat},F]$. It can equally be regarded as an even permutation of the half-diminished seventh $F^{\phi 7}=[F,A_{\flat},C_{\flat},E_{\flat}]$.
In this connection a glance at Figure 2 is revealing. One sees that Chopin’s resolution takes the initial $G_{\sharp}m$ along the path of a $2p$-decacycle to its polar opposite $B_{\flat}M$. From $G_{\sharp}m$ we follow the ‘floppy bow tie’ trajectory to $BM$, then to $E_{\flat}m$, then across to $E_{\flat}M$, up to $Gm$, and finally to $B_{\flat}M$ at the opposite side of the decagon. It seems that it is arrival at the polar opposite that gives this resolution its firmness. The situation with Wagner’s ambivalent resolution is less decisive: he merely moves the chord to an adjacent one, one step clockwise along the unique 3pp-hexacycle to which both chords belong from $G_{\sharp}m$ to $EM$. The simplicity of this transformation suggests why its character is so unsettled: nothing has changed. One is moving sideways, from indecisiveness to indecisiveness. Indeed, we submit that by studying the relative positions of the two underlying triads on the tonnetz and how the one chord can be reached from the other along a trajectory of a cycle one is offered insight into the musicality of these transitions.
There is a substantial literature addressing the issue of how four-note chords might be accommodated into tonnetz-like schemes (Childs 1998, Douthett & Steinbach 1998, Cohn 2012, Tymoczko 2011b, 2012, Nuño 2021). Such schemes have variously included dominant sevenths, minor sixths and minor sevenths, as well as other chords such as diminished sevenths and French sixths.  [^18] Note that $Cm^{7}$ and $E_{\flat}M^{6}$ are the same chord modulo a permutation. What is the underlying triad, $Cm$ or $E_{\flat}M$? The same is true of the minor sixth, which can also function as a half-diminished seventh. The virtue of such chords is that they can function in either capacity: they are ambidextrous, so one needs to take into account the context in which they appear to determine their function. Two historical precedents for this well known idea are Rameau’s notion of double emploi and Gottfried Weber’s Mehrdeutigkeit.  We present our own approach to tetrachord tonnetze in Section X. Nonetheless, in the meantime we take advantage of the fact that tetradic relations can be given an illuminating partial analysis in terms of the triads underpinning them.  [^19] We draw attention to Cohn (2012) in Chapter 7, ‘Dissonance’, at the section ‘Reduction to a Triadic Subset’ (pp. 142-145); NB Cohn’s remark at p. 144, ‘Wagner often treats minor triads and their $\phi^{7}$ supersets as interchangeable in his late music.’
Figure 6: Chopin, Ballade in $G$-minor, Opus 23, excerpt. The Tristan chord appears in the left hand, as an arpeggio, fortissimo, at Wagner’s pitches, but in the enharmonically equivalent form $E_{\sharp},B,D_{\sharp},G_{\sharp}]$. The descending figure in the right hand touches the same notes and is mirrored by the violins in Tristan und Isolde at the climax of the prelude. Chopin’s resolution is to $B_{\flat}^{\,7}$.
A prominent example of a resolution of a Tristan-genus chord is found in the Finale ‘Adagio Lamentoso’ of Tchaikovsky’s Symphony No. 6 in B-Minor, Opus 74, ‘Pathetique’. Tchaikovsky had been studying Wagner’s works and evidently wished to make use of the novel harmonic structures implicit in this material. Tchaikovsky’s modified version of the Tristan chord, heard memorably in the strings at the outset of the Adagio Lamentoso carries the full force of Wagner’s permutation. Recall that Wagner’s chord is $F,B,E_{\flat},A_{\flat}]$, whereas Tchaikovsky, in line with the overall key of his symphony uses a $Bm^{6}$ chord, in the permutation $D,G_{\sharp},B,F_{\sharp}]$, as in Figure 7, resolving into $F_{\sharp}^{\,7}$. If we transpose this passage down three semitones to facilitate comparison with Wagner and Chopin the result is a modified Tristan chord in the form $B,F,A_{\flat},E_{\flat}]$, which resolves to $E_{\flat}^{\,7}$.
Thus, Tchaikovsky swaps the two lower voices and the two upper voices of the Tristan chord and the resulting resolution of the $G_{\sharp}m^{6}$ is more like that of Chopin than that of Wagner, only leading to $E_{\flat}^{\,7}$ rather than $B_{\flat}^{\,7}$. Hence, as in the case of Chopin, there is a sense of definition in Tchaikovsky’s resolution; but there is no joy in it – this resolution has more the character of a grim facing up to reality taken with reluctant finality. Of course, we are speaking of the sense of emotion that Tchaikovsky creates so successfully in the opening measures of the last movement. But how is this achieved? It is by means of a $2p$-hexacycle.
Specifically, beginning at $G_{\sharp}m$ in the guise of a permuted $G_{\sharp}m^{\,6}$ (we stick with Wagner’s pitch classes), we find there is a ‘straight bow tie’ $2p$-hexacycle that takes one to $A_{\flat}M$ then up to $Cm$ then finally to $E_{\flat}M$. As in Chopin’s resolution, following the hexacycle right through to the polar position contributes to the sense of completion.
However, Tchaikovsky’s resolution to $E_{\flat}M$, since it is only across a $2p$-hexacycle rather than a $2p$-decacycle, as in the case of Chopin, thus has a more perfunctory character, that of a plagal cadence, and hence its desolation. We note, in contrast, that in the $G$-minor Ballade much of the passage leading up to the Tristan chord is based, in part, on the key of $E_{\flat}M$ (see the analysis of Chopin’s Opus 23 undertaken in Cohn 2012 at pp. 96-98).
Figure 7: Tchaikovsky, Symphony No. 6 in B-Minor, Opus 74, fourth movement, opening bars. The initial $Bm^{6}$ in the permutation $D,G_{\sharp},B,F_{\sharp}]$ resolves to $F_{\sharp}^{\,7}$. Transposed down by three semitones, the initial chord takes the form of a permuted $G_{\sharp}m^{6}$, namely $B,E_{\sharp},G_{\sharp},D_{\sharp}]$ or equally $B,F,A_{\flat},E_{\flat}]$, obtained from the Tristan chord by the permutation $1,2,3,4]\to[2,1,4,3]$. The resolution is to $E_{\flat}^{\,7}$, in contrast with Chopin’s resolution to $B_{\flat}^{\,7}$ and Wagner’s to $E^{\,7}$.
One can ask, did Wagner himself make use of Tchaikovsky’s version of the Tristan chord? Was Tchaikovsky aware of this? The reply is ‘yes’ and ‘very likely’. Tchaikovsky’s chord can be found in Parsifal, early in Act 3, at the passage where the horns mysteriously announce the return of Parsifal to the now languishing Kingdom of the Holy Grail, dressed in black armour and carrying the Sacred Spear. The Parsifal motif, which in the first act took the form of a simple fanfare involving a triplet figure leading from $B_{\flat}M$ to $E_{\flat}M$ has evolved (Figure 8) into a fanfare on $B_{\flat}m^{6}$ leading to a devastating $D_{\flat}m^{6}$. It is this $D_{\flat}m^{6}$, which Wagner presents in the form $E_{\natural},B_{\flat},D_{\flat},A_{\flat}]$, that is isomorphic to the $Bm^{6}$ that Tchaikovsky presents in the form $D,G_{\sharp},B,F_{\sharp}]$ in the Adagio Lamentoso.
Wagner’s version of the chord is pitched a tone higher than that of Tchaikovsky, but lowered by an octave with the $E_{\natural}$ doubled in the bass to add darkness and weight. It does not resolve to a dominant seventh, but rather, and only tentatively, to a $G_{\flat}m^{6}$ in the form $E_{\flat},B_{\flat\flat},D_{\flat},G_{\flat}]$, which is a Tristan chord. Despite the surrounding sense of gloom and oppression, Wagner’s resolution is ultimately conventional – in keeping with Parsifal’s character – i.e., up a fourth from $D_{\flat}m^{6}$, in line with the basic Parsifal motif – and hence consoling. The Tristan chord then resolves further, with more certainty, to another Tristan chord a tone lower given by an $Em^{6}$ in the form $C_{\sharp},G,B,E]$.
Figure 8: Wagner, Parsifal, excerpt from Act 3. The Parsifal leitmotif, which at Parsifal’s entry in Act 1 is a simple fanfare leading from $B_{\flat}M$ to $E_{\flat}M$, has evolved at Parsifal’s Act 3 entry into $B_{\flat}m^{6}$ leading to $D_{\flat}m^{6}$, which Wagner presents in the permutation $\{E_{\natural},B_{\flat},D_{\flat},A_{\flat}\}$, isomorphic to Tchaikovsky’s $Bm^{6}$ in the form $\{D,G_{\sharp},B,F_{\sharp}\}$ appearing at the outset of the Adagio Lamentoso.
Finally, the resolution proceeds to a diminished seventh $G,E,A_{\sharp},C_{\sharp}]$ reflecting Gurnemanz’s bewilderment at the arrival of the strange knight (‘Das ist der Brüder keiner!’). Wagner’s resolution of the ‘Tchaikovsky chord’ can perhaps be best understood in the context of the entire phrase leading from the horn fanfare in $B_{\flat}m^{6}$ to the Tchaikovsky chord at $D_{\flat}m^{6}$ on to the first Tristan resolution at $G_{\flat}m^{6}$ and then to the second Tristan resolution at $Em^{6}$. The starting point $B_{\flat}m^{6}$ and the end point $Em^{6}$ are in polar opposition on the Hamiltonian cycle of the tonnetz graph depicted in Figure 2. The Tchaikovsky chord, is exactly one quarter the way around the tonnetz, half way between the starting point and the end point. The first Tristan resolution is one black vertex further along on the cycle towards the end point, and can be viewed as a staging point before the final stage is reached at the second Tristan resolution.
This brief passage is highly symmetrical in its construction, spanning the full breadth of the tonnetz, even though with its gritty dissonances and abrupt conclusion, and only lasting a few seconds, it is has the character of a throwaway remark. It is interesting to observe that all four of the chords are linked by a 4pp-decacycle:
$\langle B_{\flat}m,G_{\flat}M,G_{\flat}m,AM,D_{\flat}m,EM,Em,GM,Gm,B_{\flat}M,B_{\flat}m\rangle.$ (8)
The diminished seventh $G,E,A_{\sharp},C_{\sharp}]$ that finishes off this sequence of four minor sixths representing Gurnemanz’s take on the unexpected guest is merely one semitone away from a $Gm^{6}$, and hence for Wagner’s dramatic purposes serves well as a point of exit from the decacycle. This view is not inconsistent with the idea that the Hamiltonian cycle circumnavigating the tonnetz is also playing a role – for not only is $Em$ opposite to $B_{\flat}m$ in the Hamiltonian cycle, one also finds that $Gm$ is opposite to $D_{\flat}m$. Hence the chords $B_{\flat}m$, $D_{\flat}m$, $Em$, GmGm are equally spaced around the perimeter of the tonnetz and the names of these chords spell out the notes of the diminished seventh on which the sequence ends.
Now, as for whether Tchaikovsky knew about Wagner’s use of the Adagio Lamentoso chord – well, we know that Tchaikovsky possessed a piano reduction of Parsifal and had played through it a number of times – so, indeed, he would have been well aware of Wagner’s use of the Tristan chord in that opera. Is it possible that Tchaikovsky got his idea elsewhere? Yes! Tchaikovsky’s chord can also be found at the strident opening of Chopin’s Scherzo in $B$-Minor, Opus 20, in the form $G,C_{\sharp},E,B]$, resolving to $A_{\sharp},C_{\sharp},E,F_{\sharp}]$, both with some doubling. Thus, we have $Em^{6}$ moving to $F_{\sharp}^{\,7}$, the first chord being in Tchaikovsky’s permutation and the second being, unusually, in the first inversion. Putting this into Wagner’s Tristan pitch levels, we have $G_{\sharp}m^{6}$ moving to $A_{\sharp}^{\,7}$, which we recognize as enharmonically equivalent to the resolution in the $G$-minor Ballade at the piú animato; though here since the dominant seventh appears in the first inversion it has a different character, demanding immediate further resolution to the tonic $D_{\sharp}m$, which in the key of the Scherzo is $Bm$.
V Pentatonic Tone Network
‘Musical keys and modulations do appear to arouse certain respondent moods or emotions. But is this, primarily, a matter of historical convention, of schooled expectation? What makes a minor third ‘sad’? Is G-minor, in the Western scale, intrinsically triste (and just what could such a statement signify?), or does its desolation stem from the use Mozart made of it in his great Quintet K 516? What of the reflexes of sensibility motivated by keys, pitch, chordal blocks in non-Western tonic systems? Is a pentatonic structure any less universal than ours?’ —George Steiner, Errata
According to Piston (1985), ‘The pentatonic scale, whose intervallic pattern corresponds to that of the black keys on the piano, is very ancient, having been used in the music of Oriental cultures perhaps longer even than the diatonic scale in the West. It is also the scale used in many European folksongs, particularly those of the British Isles.’ Examples of the use of pentatonic scales can also be found in the works of many of the well-known modern Western composers, including Bartok, Chopin, Debussy, Mahler, Ravel and Stravinsky, to name a few.  [^20] Piston/DeVoto’s characterization of the broad geographical range of the pentatonic scale may or may not be strongly research-backed. It would be interesting to know to what extent the claim there can be substantiated by recent ethnomusicological literature. Could there be an analogue of the tonnetz based on the pentatonic scale? Although some bare rudiments of the system of major and minor chords can be glimpsed in the pentatonic system, nothing like the rich dodecaphonic system of Western music exists, so it is not easy to imagine off hand what form a pentatonic tonnetz might take.
The theory of configurations suggests a way forward. If we have a system of music based on five tones, what kind of combinatorial geometry can be constructed? A hint comes from the work of Cayley (1846) who observed that five points in general position in $\mathbb{R}^{3}$ determine ten lines and ten planes that meet a generic plane in a configuration of ten points and ten lines such that three points lie on each line and three lines pass through each point (Coxeter 1950). Thus we obtain a $\{10_{3}\}$ which we can take as our pentatonic tone network.
Instead of triads, the major and minor chords are clusters of two and three notes, respectively. In pentatonic music these clusters possess a pleasing demeanor – none sound particularly distressing. If we take the pentatonic scale to be the notes C,D,E,G,AC,\,D,\,E,\,G,\,A, the ten two-note clusters are $\{CD\},\,\{CE\},\,\{CG\},\,\{CA\},\,\{DE\},\,\{DG\},\,\{DA\},\,\{EG\},\,\{EA\}$, $\{GA\}$, and the ten three-note clusters are $\{ACD\}$, $\{CDE\}$, $\{DEG\}$, $\{EGA\}$, $\{GAC\}$, $\{CDG\}$, $\{DEA\}$, $\{EGC\}$, $\{GAD\}$, $\{ACE\}$.
The pleasurable effect of such three-note clusters is well-known to Western musicians, even if they do not fit comfortably into the scheme of Western harmony except as a kind of attenuated ninth or eleventh. For example, $[ACD]$ can be viewed as an $Am^{11}$, with some tones removed, and $CDE$ can be viewed as a thinned out $CM^{9}$. An example can be found in Bach’s Toccata and Fugue in F major, at the entry of the fourth voice in the exposition of the fugue. It seems miraculous that the three upper voices converge to the tone cluster $[CDE]$ at the moment that the bass enters in the dominant on the note $C$ two octaves below. Bach makes it sound as if this near collision of the three upper lines is a happy accident.
Figure 9: The complete pentagon determined by five points C,D,E,G,AC,\,D,\,E,\,G,\,A and the ten lines pairwise joining them CD,CE,CG,CA,DE,DG,DA,EG,EA,GACD,\,CE,\,CG,\,CA,\,DE,\,DG,\,DA,\,EG,\,EA,\,GA acts as an aide mémoire for the incidence geometry of the ten lines and the ten triangles ACD,CDE,DEG,EGA,ACD,\,CDE,\,DEG,\,EGA,\, GAC,CDG,DEA,EGC,GAD,ACEGAC,\,CDG,\,DEA,\,EGC,\,GAD,\,ACE. The three lines forming the edges of a triangle are said to be incident with that triangle. Three triangles are then incident to each edge. For example, CDCD, DEDE and $CE$ are incident to $CDE$, whereas the $CDA$, CDGCDG and $CDE$ are incident to $CD$. The resulting geometry defines the Desargues configuration $\{10_{3}\}$.
Another example can be found in the Molto Adagio at the conclusion of the fugue in Bach’s Toccata and Fugue in $D$ Minor, where a permutation of the cluster $B_{\flat}CD]$ is accompanied by an $F$, leading to the cathartic minor plagal cadence at the terminus. But the occurrence of such diatonic tone clusters tends to be sporadic in the extended common practice and mostly for colour and effect rather than logic and syntax. Can one make music from clusters alone?
Cayley’s construction and the associated configuration of lines and planes can be easily remembered if one looks at the geometry of the pentacle shown in Figure 9 (cf. Mason 1977, p. 164). The pentacle includes the five lines forming the pentagon together with five further lines forming the inscribed star pentagon, making a total of ten lines altogether, and additionally there are ten triangles – five running around the edge of the pentagon, and five more forming the points of the stars. On the one hand, each triangle consists of three lines, and on the other hand each line is incident with three distinct triangles. This gives a three-to-three map between lines and planes, and hence defines an incidence geometry.
The theorem of Desargues, illustrated in Figure 10, is one of the high points of any treatment of basic projective geometry – see, e.g., that of Coxeter (1974). One of the curious features of this theorem is the fact the analogous result in three dimensions is to some extent obvious, since the two distinct planes in which the perspectival triangles reside necessarily intersect in a line containing the three pairwise intersections of the edges, whereas in two dimensions the proof is not so easily forthcoming.
The resulting scheme of incidence relations involving the ten points and the ten lines gives rise to the Desargues configuration, which can be viewed both as geometric and combinatorial. The fact that the Desargues configuration exists in the rational projective plane and that the relevant Diophantine problem has a solution hints at the deep connection this configuration admits with the theory of numbers.
Figure 10: Desargues theorem. A pair of triangles PQRPQR and $STU$ are in perspective from a point $V$ if and only if they are in perspective from a line. The line of perspective passes through the three intersection points $X=QR\cap TU$, $Y=PR\cap SU$, and $Z=PQ\cap ST$. The incidence geometry of the ten points P,Q,R,S,T,U,V,X,Y,ZP,\,Q,\,R,\,S,\,T,\,U,\,V,\,X,\,Y,\,Z and the ten lines PRY,PQZ,QRX,PRY,\,PQZ,\,QRX, SUY,STZ,TUX,PSV,QTV,RUV,XYZSUY,\,STZ,\,TUX,\,PSV,\,QTV,\,RUV,\,XYZ defines the Desargues configuration.
In fact, the musical labelling of the Desargues configuration shown in Figure 11 is in some respects even more memorable than the traditional geometric labelling of Figure 10. The five tones of the pentatonic scale yield ten two-note ‘major’ clusters and ten three-note ‘minor’ clusters. The pitches of these tones may vary from culture to culture even if the combinatorial geometry of the configuration is invariant. One could even consider a form of pentatonic equal temperament by setting $C=1$,  $D=2^{1/5}$,  $E=2^{2/5}$,  $G=2^{3/5}$,  $A=2^{4/5}$. We refer to the resulting clusters of two and three tones as the ‘chords’ of pentatonic music. Each major chord can be found in three minor chords and each minor chord contains three major chords. For example, the major chord $CD$ can be found in the three minor chords $CDE$, CDGCDG, and $CDA$, whereas the minor chord $CDE$ contains the major chords $CD$, CECE, and $DE$, as one sees in Figure 11. Here, to ease the notation we drop the braces around unordered chords. Thus we obtain a three to three map between the major chord set and the minor chord set.
Figure 12 shows on the left how the chords of the pentatonic tone network can be arranged in the form of a Levi graph, of which another view can be found on the right in which some of the other cycles can be easily discerned. What is remarkable is the overall affinity of the underlying combinatorial geometries of the pentatonic tonnetz and the Eulerian tonnetz. The results are summarized in Proposition 3. The pentatonic tonnetz can be readily used as a compositional tool. In Figure 13 we give an example of such a composition.
Looking at the Desargues configuration in Figures 10 and 11 one might be tempted to assign some special significance to the point of perspective $V$ or in the musical picture the chord $GA$. But this is illusory – any of the ten points can be taken as a point of perspective, from which the two associated triangles can be identified and the line of perspective.  [^21] Nonetheless, once the symmetry is broken by the choice of a point of perspective the other points line up to it in a specific way with certain relations among them. Thus, the choice of a preferred chord, such as $GA$, as the point of perspective is analogous to a choice of key. No one choice is preferred a priori, but once a choice has been made then the music associated with that chord involves a specific set of relations between that chord and the other chords. This is similar to the way in which in common practice tonality no specific key is preferred a priori to another – they are all on an equal footing – but once a key has been chosen then much else follows. The two-letter ‘major’ chords determine ten major keys; and since the configuration is self-dual, the ten three-letter ‘minor’ chords determine ten minor keys. A more symmetrical way of drawing the Desargues configuration can be found in figure 2.3a of Coxeter (1974) in which two different points of perspective, and hence two different lines of perspective, are treated in a manifestly symmetric way.
Figure 11: The pentatonic tonnetz of the Desargues configuration. The triangles $(EG,DG,CG)$ and $(EA,DA,CA)$ are in perspective from the point GAGA. The sides CEACEA and CEGCEG meet at the point CECE; whereas CDGCDG and CDACDA meet at CDCD; and the sides DEADEA and DEGDEG meet at $DE$. The points $CD$, CECE and $DE$ are incident with line of perspective CDECDE.
Figure 12: The pentatonic tonnetz is represented here in two isomorphic ways as the Levi graph of the Desargues configuration. Each two-note chord belongs to three distinct three-note chords and each three-note chord contains three distinct two-note chords. The girth of the graph is six, pointing to the existence of hexacycles, for example $\langle CDA,\,CD,\,CDG,\,CG,\,CGA,\,CA,\,CDA\rangle$. The graph is manifestly self-dual. One also sees that there is a perimeter Hamiltonian cycle embracing all twenty major and minor chords. In fact, there are twenty-four such icosacycles altogether. On the right, the twenty hexacycles of the pentatonic tonnetz are clearly visible.
Figure 13: A simple composition, ‘On the Perimeter’ based on a Hamiltonian cycle of the pentatonic tonnetz, using a Japanese pentatonic scale with the five notes $C$, $D$, $E_{\flat}$, $G$, $A_{\flat}$. The Hamiltonian 0pp-icosacycle is represented in this piece by the five overlapping 1pp-hexcycles of the pentatonic tonnetz, each played twice. See the Appendix, Table VII, for a tabulation of the cycles of the Desargues tonnetz.
Proposition 3.
A tonnetz can be constructed for pentatonic music in the form of a self-dual $\{10_{3}\}$ in $\mathbb{R}^{2}$ known as the Desargues configuration, in accordance with which $10$ two-note ‘major’ chords are represented by points and $10$ three-note ‘minor’ chords are represented by lines. The $30$ incidence relations between the $10$ points and the $10$ lines determine the edges of the corresponding Levi graph.
For any system of music with an odd number of tones per octave, with or without equal temperament, the combinatorics of the pentatonic system can be extended as follows:
Proposition 4.
Let a scale contain an odd number $m=2k+1$ of distinct pitches ((notes)), $k\in\mathbb{N}$. Then there are $m\choose k$ many $k$-note collections of notes, $k$-chords, that can be formed from the scale and there are $m\choose k+1$ = $m\choose k$ many $(k+1)$-note collections of notes, $(k+1)$-chords, that can be formed from the scale. Every $k$-chord is contained in k+1k+1(k+1)(k+1)-chords, and every $(k+1)$-chord contains k+1k+1kk-chords. There exists a $k+1$ to $k+1$ map between the set of $k$-chords and the set of $(k+1)$-chords, and the associated incidence structure contains no tetracycles. Hence, if we write $a$ = $m\choose k$ and $b=k+1$, the construction yields a self-dual combinatorial configuration of type $\{a_{b}\}$.
Thus, we obtain a family of tonnetze based on sets of odd cardinality. When $m=5$ we recover the pentatonic tonnetz. For a heptatonic scale ($m=7$) one obtains a configuration $\{35_{4}\}$ of 3-note and 4-note chords. For $m=9$, we get a $\{125_{5}\}$ of 4-note and 5-note chords. Even the case $k=1$ is interesting, which gives rise to a scale with three tones, together with a system of three one-tone major chords and three two-tone minor chords, for which the associated Levi graph consists of a single hexacycle.
VI Subset relations vs voice-leading relations
A possible criticism of our approach to the construction of a tonnetz for the pentatonic system (and for the more general systems of tonnetze considered in Proposition 4) might be that it is based on subset inclusion relations rather than voice-leading relations along the lines of the familiar $L$, $P$ and $R$ operations of neo-Riemannian theory that form the basis of the Eulerian tonnetz.
Prima facie, these do seem to be very different constructions. Nonetheless, it has to be admitted that the Levi graphs of the Eulerian configuration and the pentatonic configuration bear a striking ‘family resemblance’, suggesting some commonality in the superficially divergent methodologies we have employed.
Figure 14: Levi graphs for the note-to-major-chord tonnetz and the note-to-minor-chord tonnetz. On the left, each major chord (white vertex) is joined to three notes (black vertex) and each note is joined to three major chords with the obvious set-inclusion relations. The resulting graph has the same incidence structure as that of the Eulerian tonnetz, as do also the corresponding geometrical configuration and the associated hexagonal tessellation. An analogous graph can be constructed for the minor chords, on the right, again with the same incidence structure.
This suggests that we should look at subset relations in the Eulerian tonnetz. In this respect, some interesting observations can be made. Consider the twelve notes of the chromatic scale and the twelve major chords. Clearly, each chord contains three notes; but it is also the case that each note is contained in exactly three major chords. For example, the note $C$ is contained in $CM$, FMFM and $A_{\flat}M$. So, we have a three-to-three map between the note set and the major chord set. It is true as well that any note is contained in exactly three minor chords – so one obtains another three-to-three map. What do these maps look like? We can construct the relevant Levi graphs, and the results are shown in Figure 14. Curiously, each of these graphs has the incidence structure of the Eulerian tonnetz. So, evidently, the combinatorial structure of the tonnetz is already implicit in the relation between the major chords and the tones within these chords. More precisely, we have the following:
Proposition 5.
The Eulerian tonnetz can be constructed purely on the basis of set inclusion relations, without reference to voice-leading relations, minimal work or parsimony.
Proof.
Begin with the twelve notes of chromatic scale and the twelve major chords. Associated with each note are the three major chords to which it belongs. For example, the note $C$ belongs to the chords $CM$, $A_{\flat}M$ and $FM$. One can then construct a hexagonal tessellation of the plane in which each note is surrounded in a symmetric way by its three associated major chords. Each hexagon in the tessellation then consists of a cycle of notes and major chords. For example, the note $C$ is contained in the cycle $\langle C,A_{\flat}M,E_{\flat},E_{\flat}M,G,CM,C\rangle$. Now, the three notes in any given cycle can be combined to form a minor chord, so we mark that the center of that hexacycle with that minor chord. For instance, the cycle just considered is marked with $Cm$ at the center. Each minor chord is then connected to the three notes that it contains and to the three major chords in the hexagon at the center of which the given minor chord sits. The resulting triangular tessellation of the plane is the complete triadic tonnetz shown in Figure 15. If we rub out the notes and the links between notes and chords, we are left with the Eulerian tonnetz of Figure 1. ∎
Figure 15: The complete triadic tonnetz is a face-centered hexagonal tessellation admitting an infinite tripartite graph based on an underlying triangular tessellation. This infinite collection of face-centered hexagons, each containing a pitch at the center, has the property that each major chord is surrounded its three constituent pitches and each minor chord is surrounded by its three constituent pitches. If the pitches and the edges joining them are rubbed out, then one is left with the major-minor tessellation of Figure 1; and then if the enharmonic-equivalent pitches are identified, one is led to Figure 2. If one or the other of the remaining classes of edges is rubbed out instead, then by the same procedure one obtains the Levi graphs of Figure 14.
The idea that the Eulerian tonnetz can be constructed on the basis of set-inclusion relations may come as a surprise, but the matter can be understood readily enough. It is well-known that the tonnetz has two different ‘chicken-wire’ representations, one in which the notes are at the vertices and one in which the major and minor chords are at the vertices.
The note-based version of tonnetz is essentially that of Euler and has been used by Cohn (2012) and many others as a starting point for musical analyses. Chord-based versions of the tonnetz can be found in Waller (1978), Douthett & Steinbach (1998), Tymoczko (2011b) and other references. We say that a major chord and a minor chord are incident if they share two notes in common. This relation also functions as a basis for voice leading. The point is that one need not start from that relation as the primary one. Alternatively, if one takes the note/major-chord tonnetz of Figure 14 as given, then we observe that the minor chords are in one-to-one correspondence with the 2pp-hexacycles of this graph. The three minor chords to which a note belongs are the three 2pp-hexacycles that contain that note.
One sees, for example, that the note $C$ is contained in the 2pp-hexacycles corresponding to $Cm$, AmAm and $Fm$. This point of view is interesting since it breaks the duality between the major and minor chords that seems to dominate the usual philosophy of the tonnetz and to some extent restores the primacy of the major chords. According to this view, the minor chords can be identified as secondary structures – namely, the 2pp-hexacycles. In that case, one can also ask about the role of the four 3pp-hexacycles of the note/major-chord tonnetz. Inspection of the Levi graph shows that these correspond to the four augmented triads.
In Figure 15, the note set and the two chord sets are combined in the form of a face-centered hexagonal tessellation that treats all three of these sets (the set of notes, the set of major chords, and the set of minor chords) on an equal footing. A similar graph appears in Cubarsi (2024). There are three types of vertices (notes, major chords, minor chords) and three types of edges (note-major, note-minor, major-to-minor). Each note vertex is joined by three note-major edges and three note-minor edges. Each major vertex is joined by three major-note edges and three major-minor edges. Each minor edge is joined by three minor-note edges and three minor-major edges. In this way, we obtain the tripartite graph of a tripartite combinatorial configuration that can be represented, in an obvious notation, by a symbol of the form $\{12_{3,3},12_{3,3},12_{3,3}\}$.
VII Remarks on the Twelve-Tone System
The twelve-tone system developed by Arnold Schönberg, Anton Webern, Alban Berg and their followers in the Second Viennese School provided an approach to composition that was both disciplined and flexible and in the right hands capable capable of embodying a full range of musical expression. Yet, apart from the general attributes of intelligence and musicality held in common with other such works of art, the character of the music thus produced is so different from that of the preceding generations that, even if one allows for the extremes of extended tonality, we are tempted to speculate that fundamentally different structural principles are in operation, not merely tweaks of the ancien régime. Could it be that another tone network is involved?
The role of unordered hexachords in twelve-tone music has long been recognized. For example, Lewin (1968) approaches the structure of Schönberg’s Violin Fantasy with a study based on the unordered hexachord $F,G,A,B_{\flat},B,C_{\sharp}]$ and its complement $C,D,D_{\sharp},E,F_{\sharp},G_{\sharp}]$.
Unordered hexachords and their complements also played a prominent role in the work of Josef Matthias Hauer, whose techniques anticipated those of Schönberg. It is not unreasonable then to take the view that at the very outset of any approach to the theory of twelve-tone music the combinatoric geometry of such unordered sets should be investigated in the spirit we have pursued the Eulerian and pentatonic tonnetze.
Our musical starting point is a hexachord $H_{0}=\{1,2,3,4,5,6\}$ – that is, an unordered set of six distinct tones from the chromatic scale with equal temperament, whose elements we have labeled with the numbers one to six. These numbers are not, of course, the pitch classes of the six tones; their role is purely combinatoric. For example, in the first hexachord mentioned above, we set $1=F$, $2=G$, $3=A$, $4=B_{\flat}$, $5=B$, $6=C_{\sharp}$.
At first glance, such an assignment might seem structureless; but the permutation group for six objects is very special and the highly nontrivial structure of this group is directly related to the particular type of tone network that we wish to propose for twelve-tone music. The method that we consider is based on Sylvester’s theory of duads and synthemes (Sylvester 1844, 1861).
A duad is an unordered pair of distinct tones taken from $H_{0}$. A little thought shows there are fifteen duads. Thus, a duad $D$ is an unordered set of the form
$D=\{(i,j):i,j\in H_{0},\,i\neq j\}.$ (9)
By a syntheme we mean an unordered triple of duads taken in such a way that each element of $H_{0}$ appears exactly once as one of the elements of the three duads. Hence, a syntheme (Sylvester’s term, from the Greek s¡un t’ijhmi) is a set of the form
$S=\{(i,j),(k,l),(m,n):i,j,k,l,m,n\in H_{0},{\rm\,pairwise\,\,distinct}\}.$ (10)
There are fifteen synthemes altogether. Now, each syntheme contains three duads. But it is also the case that each duad is contained in three distinct synthemes. The duad $(i,j)$ is contained in three distinct synthemes, namely $\{(i,j),(k,l),(m,n)\}$, $\{(i,j),(k,m),(n,l)\}$ and $\{(i,j),(k,n),(l,m)\}$, where $i,j,k,l,m,n$ are pairwise distinct. For example, the syntheme $\{(1,2),(3,4),(5,6)\}$ contains the duads $(1,2$), $(3,4)$, $(5,6)$; and the duad $(1,2)$ is contained in the synthemes $\{(1,2),(3,4),(5,6)\}$, $\{(1,2),(3,5),(4,6)\}$, $\{(1,2),(3,6),(4,5)\}$.
Thus, we have constructed a combinatoric configuration. More precisely, we have two sets, viz., the set of duads and the set of synthemes. Each has fifteen elements and they are in three-to-three correspondence. We say that the duad $D$ is incident with the syntheme $S$ if it is among the three duads contained in $S$. Inspection of Tables II and IV shows that two synthemes have at most one duad in common. This gives us an incidence structure and we can construct the corresponding Levi graph.
Letters as totals of number-synthemes
$a$ 12, 34, 56 13, 25, 46 14, 26, 35 15, 24, 36 16, 23, 45
$b$ 12, 34, 56 16, 24, 35 15, 23, 46 13, 26, 45 14, 25, 36
$c$ 13, 25, 46 16, 24, 35 12, 36, 45 14, 23, 56 15, 26, 34
$d$ 14, 26, 35 15, 23, 46 12, 36, 45 16, 25, 34 13, 24, 56
$e$ 15, 24, 36 13, 26, 45 14, 23, 56 16, 25, 34 12, 35, 46
$f$ 16, 23, 45 14, 25, 36 15, 26, 34 13, 24, 56 12, 35, 46
Table 1: Each of the six letters corresponds to a collection of five number-synthemes called a number-total. In each number-total one finds fifteen number-duads.
Duads of letters as synthemes of numbers
abab 12, 34, 56 bcbc 16, 24, 35 cece 14, 23, 56
acac 13, 25, 46 bdbd 15, 23, 46 cfcf 15, 26, 34
adad 14, 26, 35 bebe 13, 26, 45 dede 16, 25, 34
aeae 15, 24, 36 bfbf 14, 25, 36 dfdf 13, 24, 56
afaf 16, 23, 45 cdcd 12, 36, 45 efef 12, 35, 46
Table 2: Each duad of letters corresponds to a syntheme of numbers. Each letter belongs to five distinct letter-duads and hence corresponds to a total of five synthemes.
Numbers as totals of letter-synthemes
11 $ab,cd,ef$ $ac,be,df$ $ad,bf,ce$ $ae,bd,cf$ $af,bc,de$
22 $ab,cd,ef$ $af,bd,ce$ $ae,bc,df$ $ac,bf,de$ $ad,be,cf$
33 $ab,cf,de$ $ad,bc,ef$ $ae,bf,cd$ $ac,be,df$ $af,bd,ce$
44 $af,be,cd$ $ac,bd,ef$ $ad,bf,ce$ $ae,bc,df$ $ab,cf,de$
55 $ab,ce,df$ $ae,bd,cf$ $ac,bf,de$ $ad,bc,ef$ $af,be,cd$
66 $af,bc,de$ $ad,be,cf$ $ae,bf,cd$ $ac,bd,ef$ $ab,ce,df$
Table 3: Each of the six numbers corresponds to a collection of five non-overlapping letter-synthemes called a letter-total. In each letter-total one finds fifteen letter-duads.
Duads of numbers as synthemes of letters
$12$ $ab,cd,ef$ $23$ $af,bd,ce$ $35$ $ad,bc,ef$
$13$ $ac,be,df$ $24$ $ae,bc,df$ $36$ $ae,bf,cd$
$14$ $ad,bf,ce$ $25$ $ac,bf,de$ $45$ $af,be,cd$
$15$ $ae,bd,cf$ $26$ $ad,be,cf$ $46$ $ac,bd,ef$
$16$ $af,bc,de$ $34$ $ab,cf,de$ $56$ $ab,ce,df$
Table 4: Each duad of numbers corresponds to a syntheme of letters. Each number belongs to five distinct duads and hence corresponds to a total of five letter-synthemes.
There is one more layer of structure, which comes as a surprise, and this is the following. Consider a collection of five synthemes with the property that each duad appears exactly once. This makes sense since there are fifteen duads and a collection of five synthemes must include fifteen duads. A collection of five ‘non-overlapping’ synthemes is called a ‘total’ (again, Sylvester’s term). There are exactly six such totals that can be constructed. Let us label the six totals with lower case Roman letters. Thus, the totals are elements of the set $\{a,b,c,d,e,f\}$. Since we have a second set of six elements, we can form the corresponding ‘duads’ and ‘synthemes’ for this set as well. We shall call these letter-duads and letter-synthemes, respectively. Thus, a letter-syntheme is an unordered triple of letter-duads. Then by a letter-total we mean a collection of five non-overlapping letter-synthemes. Two distinct totals have one syntheme in common. Hence, the letter duads of the totals are the original synthemes; and it follows that the letter-synthemes are the original duads. The relations between duads, synthemes, totals, letter-duads, letter-synthemes, and letter-totals can be seen in Tables I – IV (Richmond 1900, Baker 1925, Coxeter 1958).
The advantage of using the number-duads and the letter-duads is that this allows us to label the resulting Levi graph neatly, where the number-duads, which correspond to letter-synthemes, are represented by white vertices and the letter-duads, which correspond to number-synthemes, are represented by black vertices. In summary, one sees that there is indeed a rich hierarchical structure implicit in the specification of a hexachord.
The Levi graph of the Cremona-Richmond $\{15_{3}\}$ is shown in Figure 16 (Tutte 1947, Coxeter 1950, Brier & Bryant 2021). We can take this graph as representing a tonnetz for the twelve-tone system when an unordered hexachord has been selected as a primitive. Its girth is eight: there are no hexacycles.
A construction of the $\{15_{3}\}$ is as follows. Here we find it convenient to use the language of projective geometry. We begin with six points $p_{1},.\,.\,.,p_{6}$ in general position in the four-dimensional real projective space $\mathbb{RP}^{4}$. When joined pairwise, these points determine fifteen lines. Let us write $L_{12}=p_{1}\vee p_{2}$ for the line joining $p_{1}$, $p_{2}$, and $H_{3456}=\vee(p_{3}p_{4}p_{5}p_{6})$ for the complementary hyperplane joining $p_{3}$, $p_{4}$, $p_{5}$, $p_{6}$. This hyperplane is an $\mathbb{RP}^{3}$. Each of the fifteen lines meets its complementary hyperplane at a point. Thus, for example, we can write $P_{12}=L_{12}\wedge H_{3456}$ for the meet of $L_{12}$ and the hyperplane $H_{3456}$. There are evidently fifteen such ‘meeting points’. On the other hand, for example, we also have $P_{34}=L_{34}\wedge H_{1256}$ and $P_{56}=L_{56}\wedge H_{1234}$. Let us write $S_{12,34,56}$ for the line given by the meet of the three hyperplanes. Since $H_{1234}$ and $H_{1256}$ both contain $L_{12}$ it follows that the plane of intersection $D_{12}$ of these two hyperplanes contains $L_{12}$. Then since $D_{12}$ meets $H_{3456}$ at $S_{12,34,56}$ it must be that $L_{12}$ meets $S_{12,34,56}$ at $P_{12}$. By symmetry, all three of the ‘duad’ meeting points $P_{12}$, $P_{34}$, $P_{56}$ lie on the ‘syntheme’ line $S_{12,34,56}$. The fifteen meeting points are the duads of the configuration; and the fifteen lines of three-at-a-time hyperplane intersection, each of which contains three of the duads, are the synthemes of the configuration. That gives the four dimensional version of the construction, which can then be projected to a plane to provide a $\{15_{3}\}$ in two dimensions, the Cremona-Richmond configuration. Once the result has been obtained on the real projective plane, it can be Euclideanized and that takes us back to Figure 17. For further details of the construction see Richmond (1900), Baker (1925) and Coxeter (1950). Summing up the results of this section, we have the following:
Proposition 6.
A tonnetz can be associated with the selection of an unordered hexachord in twelve-tone music, given by a self-dual $\{15_{3}\}$ in $\mathbb{R}^{2}$ known as the Cremona-Richmond configuration. In this tonnetz, the $15$ ‘major’ duad chords are represented by points and the $15$ ‘minor’ syntheme chords are represented by lines. The $45$ incidence relations between the $15$ points and the $15$ lines determine the edges of the corresponding Levi graph, which is known as Tutte’s $8$-cage.
The cycles of the twelve-tone tonnetz are summarized in the Appendix, Table VIII. It is interesting to note that there are 144 Hamiltonian cycles altogether. There are no hexacycles, as we said, but there are 90 octacycles. Among these are the five overlapping 1pp-octacycles, which range about the perimeter Hamiltonian in a fashion that is very similar to the arrangement of the five overlapping 1pp-hexacycles in the pentatonic tonnetz. One might not have thought a priori that there were any particular affinities between the pentatonic system and the twelve-tone system, but there are. In fact, there is a curious kind of ‘five-ness’ that pervades the twelve-tone system that is clearly evident in our approach (e.g., in Table VIII) though not perhaps so obvious in traditional treatments of serialism. In Figure 18 we present a short composition for violin based on this pentagonal symmetry.  [^22] If any one tone of the hexachord is singled out (in this case, the tone 4), the remaining tones can be arranged with pentagonal symmetry. One example of ‘five-ness’ that does manifest itself in the traditional approach to the twelve-tone system is the involutory map from the set of integers mod 12 to itself given by multiplication by 5.
Figure 16: The Levi graph of the Cremona-Richmond configuration $15_{3}$ maps to a tonnetz for the system of twelve-tone music based on an unordered hexachord. Each white vertex represents a duad (number pair) and each black vertex represents a syntheme (letter pair). Each duad belongs to three synthemes and each syntheme contains three duads.
Figure 17: The tonnetz corresponding to an unordered hexachord maps to a configuration $\{15_{3}\}$ in $\mathbb{R}^{2}$ consisting of fifteen points and fifteen lines having the property that three points lie on each line and three lines pass through each point. The points correspond to duads or pairs of tones and the lines correspond to unordered triples of duads.
Figure 18: ‘Decacycle for Violin’, based on a 5pp-decacycle of the 12-tone tonnetz constructed from a hexachord of the form $1,2,3,4,5,6]=[F_{\sharp},G_{\sharp},C_{\sharp},D_{\sharp},E,A_{\sharp}]$. The duads and synthemes of the decacycle are given by $\langle 12,ab,56,df,13,ac,25,bf,36,cd,12\rangle$, forming the sides and vertices of a pentagon in the Cremona-Richmond configuration of Figure 17. The structure of this pentagon can also be seen in Figure 16. The relations between the letter duads and the number synthemes are shown in Table II. The note $B$ is used in this composition to ground the music.
VIII Tone Networks and Tessellations
In much of the modern literature of the Eulerian tonnetz, emphasis has been placed on the role of the 3pp-hexacycles (Cohn’s hexatonic cycles) in understanding voice leading in the analysis of music. Although the significance of this system is undeniable, one wonders if the attention given to these hexacycles has been at the expense of bypassing the consideration of other structures that are there in the mathematics and hence ought to play a role in the associated soundscape. We have given the examples of the $2p$-hexacycles, the $1p$-octacycles, the $3p$-octacycles and the $2p$-decacycles, whose position alongside the four $3p$-hexacycles might be of comparable significance. But there may be other mathematical chord system structures that are of interest as well.
To illustrate the point, we give a further example of a tone network that can be constructed by purely combinatorial arguments and yet admits a geometric representation that is not unpleasing. Here we consider again the sets of major and minor chords and we construct another map between these sets, this time of a different character.
It seems natural that given a major chord one might ask for all the minor chords that can be constructed from it by holding one note fixed and altering the other two. The spirit of this construction is similar to that of the original tonnetz. Thus (as a brief exercise in musical combinatorics will verify), given $CM$ we can construct three minor chords differing from it at two of the tones – namely, $Fm$, GmGm and $C_{\sharp}m$. Similarly, if we are given $Cm$, then we can construct three major chords differing from it at two of the tones – namely, $FM$, GMGM and $BM$.  [^23] The single common tone preserving relations between triads that we have described are known as the ‘obverse’ relations of the usual $P$, $L$ and $R$. For example, $CM\to Fm$ is $L^{\prime}$, the obverse of $L$. These terms were coined by Morris (1998).  Interestingly, under this new tonnetz the major and minor chords split into two distinct classes. More precisely, the associated bipartite graph is not connected, but rather is given by the union of two distinct connected components. Thus we obtain two sets of twelve chords, each containing six major chords and six minor chords, as shown in Figure 19. Since the presence of any one chord in one of these Archimedean tonnetze determines the rest, we can call the first one the $C$-major Archimedean tonnetz and the other one the $C$-minor Archimedean tonnetz.
Figure 19: The Archimedean tonnetze. Each major chord can be changed into a minor chord by shifting two of the tones, and each minor chord can be changed into a major chord by shifting two of the tones. The resulting bipartite graph has two connected components, the $CM$ Archimedean tonnetz (on the left) and the $Cm$ Archimedean tonnetz (on the right).
As with the other tonnetze we have considered, one can read off the resulting cycles directly from the bipartite graph (see Appendix, Table IX). There are hexacycles, octacycles, decacycles and dodecacycles, but also tetracycles such as $\langle Gm$, $F_{\sharp}M$, $C_{\sharp}m$, $CM$, $Gm\rangle$. The significance of the tetracycles is that these graphs are not the graphs of configurations – for in order for a bipartite graph to be the Levi graph of a configuration it must be of girth at least six, which has indeed been the case in the previous examples that we have considered. Nonetheless, if we lay out identical copies of the dodecagons in a lattice and connect adjacent vertices across the lattice (rather than across the dodecagon), what emerges is a tessellation of the plane, composed of dodecagons, hexagons, and squares, as in Figure 20.
Figure 20: Tessellation of the plane with dodecagons, hexagons and squares generated by the $CM$ Archimedean tonnetz. A similar tessellation can be constructed for the $Cm$ Archimedean tonnetz.
This tessellation is one of a number known to Kepler (1619). Each dodecagon is surrounded by six hexagons and six squares in a remarkable alternating pattern known as an Archimedean or semi-regular tessellation  [^24] The internal angle of two adjacent sides of a regular $n$-gon is $(n-2)\pi/n$. It follows that if we wish to fit three polygons all with sides of the same length around the same vertex, then the numbers of sides $n_{1},n_{2},n_{3}$ of the polygons must be so that $(n_{1}-2)\pi/n_{1}+(n_{2}-2)\pi/n_{2}+(n_{3}-2)\pi/n_{3}=2\pi$. This condition is satisfied by $n_{1}=4$, $n_{2}=6$, $n_{3}=12$. More generally, there are 21 different solutions with various numbers of polygons and various numbers of sides to fit a collection of regular polygons around a vertex. Of these, 11 can be extended across the plane, giving the 11 Archimedean tilings. of type $\{4,6,12\}$. See, for example, Grünbaum & Shephard (1977, 2016), Conway, Burgiel & Goodman-Strauss (2008) and Wilson (2016).
Since the chords at adjacent vertices share one tone in common, we can associate each edge with a tone. The three edges meeting a vertex give the tones of which the chord corresponding to that vertex is composed. The Hamiltonian cycles in the Archimedean tonnetze ascend in the circle of fifths going counterclockwise around the rim. As a result, two of the three chords a given chord is adjacent to are a fourth or fifth apart.  [^25] A dual tessellation called the ‘Laves tiling’ associated with $\{4,6,12\}$ can be constructed (Grünbaum & Shephard 2016, at pp. 95-98) with the property that the vertices of the original tiling correspond to the tiles of the dual tiling and the tiles of the original tiling correspond to vertices of the dual tiling. The edges of the original tiling and its dual can then be identified. The Laves tilings associated with the Archimedean tonnetze can be seen in Rietsch (2024) at pp. 336-338. Each dodecagon of the original tonnetz corresponds to a vertex of the corresponding Laves tessellation met by twelve edges. Thus, our geometric/combinatorial method gives a construction of the duals of Rietsch’s $G_{2}$ tonnetze, which were obtained with Lie theory.
Figure 21: A simple ‘Fanfare for Kepler’ based on a hexacycle of the $CM$ Archimedean tonnetz. The hexacycle $\langle CM,C_{\sharp}m,A_{\flat}M,Am,EM,Fm,CM\rangle$ is a closed sequence of major and minor chords with the property that each chord differs from its successor at exactly two tones.
The Archimedean tonnetze can be used as a basis for composition, and they are rather useful.
The fact that two notes of a triad change at each transition ensures that the chord sequences are less bland or trance-like than the cycles of the Eulerian tonnetz, and hence are more interesting to listen to even in the absence of the rich tonal environment from which historical examples of the use of the Eulerian tonnetz are usually extracted.  [^26] According to Piston (1985), ‘The purpose of this procedure [the practice of voice leading in harmonic progression] is to insure the smoothest possible connection of two chords, so that one seems to flow into the next. Continued practice of the process will, however, result in rather dull music.’   In Figure 21 we illustrate this point with a short composition in the form of a fanfare based a cycle of the $CM$ Archimedean tonnetz. The occasional common practice modalities embedded in this tonnetz, like the plagal cadence from $Fm$ to $CM$, help to moderate the effects of the more unusual chord transitions and ground the piece with an aura of familiar tonality.  [^27] It is interesting to note how different cycles of the Archimeden tonnetz have contrasting voice-leading properties, along the lines of Tomoczko (2020), pp. 116-128. A counterclockwise movement through the tetracycle $\langle CM,C_{\sharp}m,F_{\sharp}M,Gm,CM\rangle$ induces a one-step scalar transition, whereas movement through a hexacycle returns each voice to its starting pitch, and a clockwise trajectory through the 0pp-dodecacycle leads every voice up an octave. Thus, the figure reflects aspects of the topology of voice-leading spaces. In this respect the tessellation in Figure 20 has an advantage over the associated bipartite graph of Figure 19 since the various cycles are visually more distinct.
IX Dialogo sopra le reti tonali
The scene is set in central London. Salviati and Sagredo sit in the Fidelio Cafe at a table for three, waiting for their friend Simplicio to join them for a drink before the concert. More and more people arrive and there is a buzz in the atmosphere in anticipation of the evening’s performance. Salviati looks at his telephone.
SALVIATI – Where is Simplicio? He always seems to be running late!
SAGREDO – Don’t worry, he’ll arrive in time for the concert. In any case, while we have our drink I would like to ask you one or two questions about that theory you were telling me about at our last meeting – what were the names of the authors?
SALVIATI – Bolando and Ugopoli. Yes, I can tell you a bit more about it if you wish.
SAGREDO – Very good. You might as well remind of what you had already said. The idea of treating certain musical structures as mathematical objects attracted my attention.
SALVIATI – Well, it’s based on the tonnetz. As you say, they take a mathematical approach. Their central argument, if I interpret it correctly, is that mathematics is not merely a kind of tool box that a music theorist might use for whatever purpose he finds it helpful. The latter would be in the tradition of old-fashioned applied mathematics, but that is not what Bolando and Ugopoli are up to. They treat music and mathematics as equal partners. They look for certain common structures – of a combinatoric, geometric or number theoretic nature – that arise naturally in both music and mathematics. So yes, they are using mathematics to explain certain phenomena in music – but they are also looking for ways in which mathematical structures manifest themselves in music – or how the same underlying idea can work its way out both in music and in mathematics.
SAGREDO – This reminds me of the way physicists use geometry. It is not simply that geometry can be used as a tool for analyzing physical phenomena, though it can be used like that, but rather that the physical phenomena are fundamentally geometric in their nature. For example, it is not merely the case that space-time can be modelled as a four-dimensional manifold with an indefinite metric – rather that space-time really is such a manifold.
SALVIATI – Yes, it’s just like that. So, likewise, they argue, in our considerations of the theory of music, we need to regard the music and the mathematics as a kind of unity. Bolando and Ugopoli begin with Waller’s representation of the tonnetz as a bipartite graph. Waller did not know about the tonnetz, and he had no particular interest in voice leading and all that, as far as I am aware. He approached the problem of the relationship between major and minor chords from a purely combinatorial and group theoretical point of view.
SAGREDO – Yes, that is what I understand. He produced that diagram of the relations between major and minor chords in the form of a regular bipartite graph. And I recall what you were saying earlier – that Bolando and Ugopoli recognized (and seem to have been the first to do so) that the Waller construction is a Levi graph – the graph of a configuration.
SALVIATI – Yes, that is exactly the point. There is an isomorphism between such graphs and configurations. These configurations are beautiful, interesting mathematical objects.
SAGREDO – I get it. What Bolando and Ugopoli are suggesting is that for every configuration there exists an abstract music – that the points and lines of the configuration represent different types of chords, analogous to major and minor, that the adjacency of a point and line corresponds to what music theorists call efficient voice leading between the chords.
SALVIATI – Correct. Bolando and Ugopoli look at various examples of abstract music systems. First they look at the usual tonnetz. They show that its Levi graph is isomorphic to a configuration of twelve lines and twelve points, what we call a $\{12_{3}\}$, with the property that three lines go through each point and three points lie on each line. Such configurations have been investigated since the 19th century and it is known that there are 229 of them. Bolando and Ugopoli were able to identify the tonnetz as the D222. This configuration has been studied a lot over the last few decades by mathematicians, so it is remarkable that no one had previously realized that it is Euler’s tonnetz.
SAGREDO – You say Euler, but I understand that Euler himself only had a sketchy idea of the picture that we now regard as the Tonnetz.
SALVIATI – True, and the idea was developed further by the visionary Hugo Riemann, but it was not until the 1980s and 1990s that the theory of the tonnetz took its modern form.
SAGREDO – Well, so I gather. But it is always in the form of those triangular patterns that one sees in Cohn’s book and elsewhere. I can see why the musicologists like them, because they can trace progressions from triangle to triangle, and understand certain aspects of harmony that were previous obscure, which is in fact quite interesting – but it seems a bit vacuous mathematically speaking, if I may say so, though maybe I’m missing something.
SALVIATI – No, you are correct, that is indeed the problem. Bolando and Ugopoli seem to be the first to have proposed the systematic use of the Levi graph of the tonnetz as a basis for musical analysis. The point is that the cycles of the Tonnetz are all clearly visible in the Levi graph. And progressions can be viewed as movement from vertex to vertex.
SAGREDO – Or even better, one can visualize progressions as movements from point to point in Euclidean space along the straight lines that are joining them.
SALVIATI – Yes, that is correct.
SAGREDO – That reminds me, I am puzzled about one thing. When I attended geometry lectures while I was doing my mathematics degree in Padova, we were taught that configurations are constructions in projective geometry, which was considered to be a rather advanced topic – but here we are talking about Euclidean geometry. Does that make any difference?
SALVIATI – Well, yes and no. The theory of configurations goes back many centuries and it wasn’t until comparatively recently, the last two hundred years or so, that the subtleties of these different types of geometry began to be fully appreciated. It is a mistake to try to pigeonhole everything and put it all in neat boxes. That is not the way mathematics works. The theory of configurations spans a lot of material and one has to take it as it is. Let me be more specific. For some purposes – and Grünbaum takes this point of view – it can be useful to confine one’s attentions exclusively to configurations of points and lines in $\mathbb{R}^{2}$.
SAGREDO (interrupting) – Isn’t that a bit restrictive?
SALVIATI – Not really – even this seemingly narrow definition embraces a wealth of material. For other purposes, one looks at configurations in projective spaces such as $\mathbb{RP}^{2}$ or $\mathbb{CP}^{2}$. There are also certain special projective spaces composed of a finite number of points and these too admit configurations – some of which may be musically interesting. In fact, one can drop the requirement that the configuration is representable over a field and look at purely combinatorial configurations. All of these configurations have Levi graphs associated with them and each defines a species of abstract music.
SAGREDO – This is making me dizzy – I think I need another drink. Where is Simplicio? Are you saying to me that each and every one of these configurations gives rise to a form of music?
SALVIATI – Well, yes – at least in principle.
SAGREDO – So we can make a list of configurations, and work our way through them.
SALVIATI – Yes, let’s do it. First, there is the $\{7_{3}\}$ of Fano, the Levi graph of which has 14 vertices, 7 white and 7 black. It does not admit a representation over the reals or the complexes, but does admit a geometric representation over the finite field $\mathbb{F}(2)$.
SAGREDO – That is the binary field.
SALVIATI – Correct. Now, where was I? Hmmm – next we have the Möbius $\{8_{3}\}$, which admits a geometric representation in $\mathbb{CP}^{2}$. And then we have the three $\{9_{3}\}$s, each of which can be realized in $\mathbb{R}^{2}$. The Pappus configuration is among these.
SAGREDO – What about the $\{10_{3}\}$ configurations?
SALVIATI – There are ten of them. One is the Desargues configuration, which Bolando and Ugopoli have used to construct a tonnetz for pentatonic music.
SAGREDO – That does not mean that existing pentatonic music always conforms to the structure of this tonnetz, merely that one can compose new pieces of music by use of it?
SALVIATI – Exactly, their pentatonic tonnetz can be used to create new music. Similarly in the case of their theory of the twelve-tone $\{15_{3}\}$ tonnetz, which amounts to a proposal for a new way that quasi-serial music might be structured.
SAGREDO – I see, though it sounds rather difficult.
SALVIATI – In fact, Bolando and Ugopoli provide what I think amounts to an entire program for the future of music. Each configuration gives rise to a form of music.
SAGREDO – Very intriguing, I must say.
SALVIATI – Now, Sagredo, where were we in our list? Well, coming back to the $\{10_{3}\}$ configurations, there is one very curious detail that I meant to mention. All but one can be realized over the reals and even over the rationals. But one of them admits no such realization. Nonetheless, it is perfectly well-defined as a combinatorial configuration and it has a Levi graph. There is nothing mysterious about it.
SAGREDO – So, it still gives rise to a music of some sort?
SALVIATI – Exactly. So, one has to beware of jumping to conclusions about mathematical objects and their properties simply on the basis of the names given to them. Bolando and Ugopoli never claimed that realizability is essential for a configuration to be musically viable. Quite the contrary. The point, rather, is that with that one peculiar exception, all the $\{7_{3}\}$, $\{8_{3}\}$, $\{9_{3}\}$, $\{10_{3}\}$, $\{11_{3}\}$ and $\{12_{3}\}$ configurations are realizable.
SAGREDO – So, all but one of the 270 tonnetze that can be constructed for scales of twelve notes or less are geometrically realizable over some field, and all but three of them are realizable in two-dimensional Euclidean space?
SALVIATI – Precisely. Nearly all of the ones that you will tend to look at in practice, at least in the first instance, fall into the ‘geometrically realizable’ category. It is not as though one is favouring or cherry-picking them in some way. They just happen to be in a heavy majority. But you can make music with any of these 270 tonnetze, whether or not they are geometrically realizable.
SAGREDO – I see. Well, then, that leads me to an important question. How do you know whether a given configuration will give rise to an interesting form of music?
SALVIATI – Well, you don’t, in advance. But I think that Coxeter’s remark – ‘interesting configurations are represented by interesting graphs’ – may be relevant here. The point is that a tonnetz should be regarded as a representation of a configuration. It follows, therefore, we may presume, that interesting configurations are represented by interesting tonnetze. This is the motivation for studying configurations as a basis for musical analysis. Interesting configurations give rise to interesting musical structures.
SAGREDO – Well, Salviati, that is a pretty sweeping conclusion.
SALVIATI – Precisely. That is exactly what our two Academician friends are saying: interesting configurations give rise to interesting musical structures – in particular, to interesting tonnetze.
SAGREDO – Hmmm, fascinating. And what about voice leading. The musicologists make lot of fuss over it, but you have scarcely mentioned it.
SALVIATI – Precisely. Well, Bolando and Ugopoli say that voice leading is a concept that is difficult to define in the abstract. In any case, music gets a bit tedious if the parts only move around in semitones or whole tones, or if the chord transitions are all like that.
SAGREDO – Of course, much has been said on this. Think of songs like The Star-Spangled Banner, or Somewhere Over the Rainbow. They are full of leaps and arpeggiated chords.
SALVIATI – True, but they also have extended passages of tonal and semitonal voice leading. The composer has to get the balance right – that is a matter of art, not science.
SAGREDO – So it’s really a question of understanding what is going on in those isolated passages where the music seems to be flowing continuously.
SALVIATI – Yes, music involves piecewise continuous processes.
SAGREDO – But what do Bolando and Ugopoli say about this?
SALVIATI – They argue that the structure of the tonnetz is implicit in the relation between major chords and the notes that make up the chords. No mention of the minor chords is needed. One eventually recovers the usual tonnetz relations between major and minor chords, but these are derived. No a priori assumptions about voice leading are required.
SAGREDO – That is intellectually satisfying, I suppose. I never bought into all that voice leading dogma. Isn’t it really a secondary matter, at least mathematically speaking?
SALVIATI – Yes, well, you can say that. You still have your voice leading in the Bolando-Ugopoli system, but it is implicit in the structure of the incidence relations of the points and lines in the configuration that defines the genre of music in which your piece is composed.
SAGREDO Ah, now that makes more sense. By the way, perhaps we should order some more wine, while we wait for Simplicio. I think we need another drink to help us along with these musico-mathematical thoughts.
SALVIATI – Agreed!
X Tone Networks for Tetrachords
‘. . . indeed, the Tristan prelude (and the opera as a whole) could be said to be “about” the various ways of resolving a Tristan chord to the dominant seventh chord . . .’
—D. Tymoczko, A Geometry of Music
We promised that we would return to the construction of tonnetze for tetrachords. The work that has been hitherto carried out in this area offers many insights into the relation of the families of tetrachords to one another but at the same time forces one to realize the daunting nature of the task. In Tristan und Isolde, the unordered minor sixth (or half-diminished seventh) chord (i.e., the Tristan chord or one of its permutations) resolves in no less than eight distinct ways into dominant seventh chords, according to Tymoczko (2012) at p. 302. If we transpose the initial chord (as does Tymoczko) to $G_{\sharp}m^{6}$, the eight listed resolutions are to the following dominant sevenths: $E^{7},F^{7},G^{7},{A_{\flat}}^{\!7},{B_{\flat}}^{\!7},B^{7},{D_{\flat}}^{\!7},{D}^{7}$. This list includes Chopin’s resolution to ${B_{\flat}}^{\!7}$ in the $G$-minor Ballade, discussed earlier, but not Tchaikovsky’s resolution to ${E_{\flat}}^{\!7}$ in the Adagio Lamentoso, which nonetheless seems to work quite well as a resolution. Perhaps it too ought to be added to the list on an honorary basis? On that argument it is tempting to believe that almost any resolution to a dominant seventh is acceptable. Is there any logic to Wagner’s choice of admissible resolutions?
Our approach to the matter of the Tristan resolutions will be to adopt a variant of the strategy that we used to create the Archimedean tonnetze. Thus, we consider all transitions from minor sixth chords to dominant sevenths with the property that exactly two tones change. The goal is then to derive a Levi graph for the resulting relations. The problem that immediately arises is that there are too many of such transitions, some of them even admitting tetracycles, such as $\langle{A_{\flat}m}^{6},$ $B^{7},$ ${Dm}^{6},$ $F^{7},$ ${A_{\flat}m}^{6}\rangle$, hence blocking the possibility of a Levi graph. Thus, we need some principle that can be applied uniformly across the class of chords being considered to limit the range of transitions sufficiently to ensure that we get a Levi graph. Our criterion for success will be that the resulting Levi graph is associated with an interesting configuration.
Consider then the following strategy. Classical music theorists have observed that in both the minor sixth and in the dominant seventh there is single dissonant tone that combines with a consonant triad. For example, the dissonant note in $Cm^{6}$ is $A$ and in $C^{7}$ it is $B_{\flat}$.
Figure 22: A Tristan-genus tonnetz. The twelve minor sixth chords and twelve dominant seventh chords can be arranged in a Levi graph with the incidence geometry of a configuration $\{12_{3}\}$. Each minor sixth chord is connected to three dominant seventh chords and each dominant seventh is connected to three minor sixths. At each transformation two notes are retained in the new chord, one being the ‘dissonant’ note of the original tetrachord, which may or may not be a dissonance in the new one. Edges that cross the tonnetz preserve the dissonant note of the tetrad as a dissonance in the new tetrad. For example, the dissonant $F$ of $G_{\sharp}m^{6}$ is mapped to the dissonant $F$ of $G^{7}$.
Therefore, let us impose the condition that the dissonant note is unchanged. That implies that two members of the underlying consonant triad have to shift, and we are back to a situation similar to what we considered earlier in connection with the Archimedean tonnetze. The result is constrained by the requirement that a minor sixth must transform into a dominant seventh and vice versa.
It is easy to see that there are exactly three transitions that send $Cm^{6}$ to a dominant seventh chord in such a way that the $A$ is retained, namely $Cm^{6}\to A^{7}$,  $Cm^{6}\to B^{7}$, and $Cm^{6}\to D^{7}$. Likewise, there are exactly three ways in which $C^{7}$ can be sent to a minor sixth such that the $B_{\flat}$ is retained, namely $C^{7}\to B_{\flat}m^{6}$,  $C^{7}\to E_{\flat}m^{6}$, and $C^{7}\to C_{\sharp}m^{6}$. Similar transformations exist by transposition for the other minor sixths and dominant sevenths. Putting these together, we obtain a connected bipartite graph of degree three and girth six for the minor sixths and dominant sevenths, shown in Figure 22. Remarkably, we thus obtain a Levi graph and a self-dual configuration $\{12_{3}\}$.
By counting cycles, one easily determines that this graph is different from the Levi graph of the Eulerian tonnetz. Now, the Eulerian tonnetz has a total of sixteen hexacycles – that is, the four ‘hexatonic’ 3pp-hexacycles and the twelve ‘pitch-retaining’ 2pp-hexacycles. But the Tristan graph in Figure 22 admits twenty-four hexacycles – specifically, we have twelve 1pp-hexacycles (hexabeanies) of the form $\langle C^{7},E_{\flat}m^{6},F^{7},G_{\sharp}m^{6},A_{\sharp}^{7},C_{\sharp}m^{6},C^{7}\rangle$ and twelve 2pp-hexacycles (bow ties) of the form $\langle C^{7},E_{\flat}m^{6},F^{7},F_{\sharp}m^{6},D_{\sharp}^{\,7},C_{\sharp}m^{6},C^{7}\rangle$.
Note that we have retained the ‘$p$’ notation for edges that cross the graph, where in the present context one can take it to mean ‘preserve’ (as in ‘preserving the dissonant note’) rather than ‘parallel’. Since our new Levi graph has a different number of hexacycles than that of the Eulerian tonnetz, this means that the two are not isomorphic. On the other hand, since it is known that all the combinatorial $\{12_{3}\}$ configurations are realizable in the Euclidean plane, one can ask which of the 229 it turns out to be, if it is not the D222. We have been able to identify it as the D228 of Daublebsky von Sterneck (1895). This rather attractive and highly symmetric configuration is plotted in Figure 23. Since they are isomorphic, we can refer to either of the Levi graph in Figure 22 and the configuration shown in Figure 23 as the ‘Tristan tonnetz’.   [^28] Using a metaphysical analogy, one can think of these three abstractions – tonnetz, Levi graph and configuration – as forming a Trinity: with the Configuration as Father, the Graph as Son, and the Tonnetz as Geist. Thus, Father, Son and Geist are distinct and yet inseparable.
Proposition 7.
A tonnetz for the Tristan genus of dominant seventh chords and minor sixth chords can be constructed by limiting the maps between these chord sets to two-note transitions that preserve the dissonant note of the chord being mapped. The resulting incidence geometry, which is distinct from that of the Eulerian tonnetz, is the self-dual configuration $\{12_{3}\}$ in $\mathbb{R}^{2}$ known as the D228 of Daublebsky von Sterneck. The $12$ dominant sevenths are represented by points and the $12$ minor sixths are represented by lines. The $36$ incidence relations between the $12$ points and $12$ lines determine the edges of the associated Levi graph, where the white vertices are dominant sevenths and the black vertices are minor sixths.
Figure 23: The geometric configuration of the Tristan-genus tonnetz is the D228 of Daublebsky von Sterneck, for which the corresponding Levi graph appears in Figure 22. This self-dual configuration is one of the four ‘homogeneous’ cases arising among of the 229 configurations of type $\{12_{3}\}$. The group of the configuration is generated by quarter-turn physical rotations of the overall figure and cyclic permutations of the three mutually inscribed squares among each other. The three squares correspond to the three $4p$-octacycles of the Levi graph.
The symmetry group of this Tristan-genus configuration is generated by quarter-turn physical rotations of the overall figure and cyclic permutations of the three mutually inscribed squares among each other. The three squares, which are on an equal footing, correspond to the three $4p$-octacycles of the Levi graph. The musical significance of these three mutually inscribed quadrilaterals of the configuration is that movements within these structures give rise some of the characteristic sonorities of late Wagner operas.
We shall look at several examples of the use of these ‘Wagnerian octacycles’. First, we return to the prelude to Act I of Tristan und Isolde. Three different versions the opening motif shown in Figure 5 appear in succession at the outset of the prelude. The first occurrence is as in Figure 5. In the second occurrence, the entire motif is transposed up by a minor third, apart from the first note, which is transposed up only one tone, so the interval between the first and the second note is a major sixth rather than the minor sixth one hears in the first occurrence of the motif. In the third occurrence, the initial rise is again by an interval of a major sixth, but the resolution of the minor sixth chord to a dominant seventh is different from that of the first two occurrences. In more detail, the three resolutions are as follows:
$(i)\hskip 7.11317ptG_{\sharp}m^{6}\to E^{7},\hskip 14.22636pt(ii)\hskip 7.11317ptBm^{6}\to G^{7},\hskip 14.22636pt(iii)\hskip 7.11317ptFm^{6}\to B^{7}.$ (11)
In an octacycle with the structure $\langle 1,2,3,4,5,6,7,8,1\rangle$, where the odd numbers represent minor sixths and the even numbers represent dominant sevenths, we note that the polar of any chord is necessarily a chord of the same mode. Thus, the polar of chord 1 is chord 5, which is half way around the cycle. But for some purposes we may be interested in the ‘most opposite’ chords of the opposite mode. We can call these the ‘subpolars’; there are two of them. Thus, the subpolars of chord 1 are chords 4 and 6. With this terminology we are ready to look at relations of the chords in the Tristan prelude. In the tables of cycle numbers shown in the Appendix we see that the Tristan tonnetz admits twelve 3pp-octacycles and three 4pp-octacycles. We observe that the polar of $G_{\sharp}m^{6}$ along the unique 4pp-octacycle containing that chord is $Dm^{6}$. The subpolars of $G_{\sharp}m^{6}$ are $E^{7}$ and $B^{7}$.
Thus, Wagner’s well known first resolution (i) runs along the octacycle beginning at $G_{\sharp}m^{6}$ and resolves at the subpolar $E^{7}$. A glance at Figure 22 shows that resolution (ii) begins on the same octacycle at $Bm^{6}$ and resolves to its subpolar, which is $G^{7}$.
Resolution (iii) again begins on the same octacycle, this time at $Fm^{6}$, but resolves to $B^{7}$, which is not on the same octacycle – but this $B^{7}$ nevertheless stands in a definite relation to $Fm^{6}$: in fact, $B^{7}$ is polar to $Fm^{6}$ along the 2pp-hexacycle (a bow tie) that connects these two chords.
In summary, all three minor sixths – $G_{\sharp}m^{6}$, $Bm^{6}$ and $Fm^{6}$ – lie on the same octacycle of the Tristan tonnetz, along with two of the resolvent dominant sevenths, $E^{7}$ and $G^{7}$. But the final resolution starting at $Fm^{6}$ is deflected along one of the three 2pp-hexacycles containing that chord, namely the one that passes through $E^{7}$ and $Dm^{6}$, to reach its polar $B^{7}$, which we recall is the other subpolar of the original $G_{\sharp}m^{6}$. Hence, all of the chords of this profound opening gesture of the Tristan prelude stand in a definite and rather subtle set of relations to one another on the Tristan tonnetz.  [^29] This passage at the beginning of the opera seems to have sprung fully formed from the composer’s mind. One is reminded of the remarkable analysis of William Blake’s Infant Sorrow carried out by the linguist Roman Jakobson (1970), who after similar consideration of the intricate syntactic and morphological symmetries of Blake’s short poem comes to the conclusion that it couldn’t have been assembled piece by piece.
Another example of the use of octacycles can be found in Act 2, Scene 2 of Tristan und Isolde – the ‘Liebesnacht’. The duet is marked by two warnings from Brangäne (‘Habet acht!’) that night will soon give way to day, following which first Tristan and then Isolde plead for death (‘Lass mich sterben!’). This section is accompanied by a well-known sequence of six three-chord progressions, each extending from minor sixth to minor sixth to dominant seventh, visiting all twelve of the minor sixth chords and half of the dominant seventh chords along the way. Let’s see what this looks like on the Tristan tonnetz. The six progressions, in two groups of three, separated in pitch by a semitone, are as follows:
$(i)\hskip 7.11317ptA_{\flat}m^{6}\to Dm^{6}\to C^{7},\hskip 14.22636pt(ii)\hskip 7.11317ptB_{\flat}m^{6}\to Em^{6}\to D^{7},\hskip 14.22636pt(iii)\hskip 7.11317ptCm^{6}\to F_{\sharp}m^{6}\to E^{7},\hskip 14.22636pt$ (12)
$(iv)\hskip 7.11317ptAm^{6}\to E_{\flat}m^{6}\to C_{\sharp}^{7},\hskip 14.22636pt(v)\hskip 7.11317ptBm^{6}\to Fm^{6}\to D_{\sharp}^{7},\hskip 14.22636pt(vi)\hskip 7.11317ptC_{\sharp}m^{6}\to Gm^{6}\to F^{7}.\hskip 14.22636pt$ (13)
In each progression the two minor sixth chords are polar opposites on the Hamiltonian cycle of the Tristan tonnetz (i.e., twelve edges apart around the rim of the tonnetz) and are also polar opposites on the unique 4pp-octacycle containing them (four edges apart). For instance, $G_{\sharp}m^{6}$ and $Dm^{6}$ are polar on the perimeter Hamiltonian cycle and on the 4pp-octacycle
$\langle A_{\flat}m^{6},G^{7},Fm^{6},E^{7},Dm^{6},C_{\sharp}^{\,7},Bm^{6},A_{\sharp}^{\,7},A_{\flat}m^{6}\rangle.$ (14)
Moreover, again in each of the six progressions, the first minor sixth chord and the resulting dominant seventh chord are simultaneously polar opposites (three edges apart) on the $2p$-hexacycle containing them and on the two $1p$-hexacycles containing them.
Next, we return to Tymoczko’s list of the eight dominant seventh chords admitted in the opera Tristan. Can we make sense of this list in the light of what we have learned? First, we observe that the names of these chords spell out the notes of the octatonic scale $O_{23}$. We note that the list splits into two sets, differing by a semitone: namely, $E^{7},G^{7},{B_{\flat}}^{\!7},{D_{\flat}}^{\!7}$, and ${D}^{7},F^{7},{A_{\flat}}^{\!7},B^{7}$. But these are the seventh chords that belong to precisely two of the three 4pp-octacycles of the Tristan tonnetz. So now we understand which chords are missing from the list: they are the four dominant sevenths belonging to the excluded octacycle.
While we are on the topic of Tristan und Isolde, let us recall the resolutions used by Tchaikovsky and Chopin that we had tentatively discussed in Section IV. How do these resolutions square up on the Tristan tonnetz? In the case of Tchaikovsky, we observe that this resolution from $G_{\sharp}m^{6}$ to $D_{\sharp}^{7}$ (in Wagner’s pitches) at the beginning of the Adagio Lamentoso is polar along the unique 1pp-hexacycle connecting these chords. In the case of Chopin’s resolution from $G_{\sharp}m^{6}$ to $B\flat^{7}$ at the transition to the più animato section of the Ballade in $G$-minor, we find that this $B\flat^{7}$ is adjacent to $G_{\sharp}m^{6}$ on the Tristan tonnetz. So indeed there seems to be a kind of tonnetz logic to Chopin’s choice that eludes classical analysis.  [^30] We have approached these resolutions both from a triadic perspective and now from a tetrachordal perspective. Why both? We are working at the boundary of art and science. It is of the nature of works of art that one can never quite pin them down. On the one hand, they resist analysis, and on the other hand they will readily embrace multiple modes of analysis. Each such approach reveals yet another aspect of the artwork. One might even say that if the object under consideration yields completely to any one mode of analysis, then it isn’t art. In this respect, we seem to be in agreement with Tomoczko (2011) at p. 304: ‘ .  .  .  I am not proposing a simple method or rule for doing musical analysis: there is no royal road to musical understanding, geometrical or otherwise.’ We enter not the debate on whether psychological notions can be meaningfully incorporated into music theory in a mathematically systematic way. All such endeavors remain speculative, though in our view certainly worth attempting. Nonetheless, even the idea of the perceived ‘distance’ between chords is fraught with difficulty. Tomoczko (2011) offers a critique of tonnetz-based distance measures at Appendix C.
We turn to consider another Wagnerian example of the resolution of minor sixths to dominant sevenths, this time in an octacycle of the Tristan tonnetz from Götterdämmerung. In the final section of the opera, at the beginning of the scene where Brünnhilde orders the vassals to pile up logs to create a great funeral pyre around the dead Siegfried’s body, there is an extraordinary well-known sequence of chords of the Tristan genus. The sequence is as follows, extending over a total of twelve measures, beginning softly at the $A^{7}$, building up to a climax at the $F_{\sharp}M$:
$A^{7}\to B_{\flat}m^{6}\to C^{7}\to D_{\flat}m^{6}\to E_{\flat}^{\,7}\to Em^{6}\to C_{\sharp}^{\,7}\to F_{\sharp}M.$ (15)
The final transition $C_{\sharp}^{\,7}\to F_{\sharp}M$ can be heard as an authentic cadence designed as a ‘wrap up’. What interests us is the sequence of six chords leading up to this. Looking at the Tristan tonnetz, one sees that all six tetrads of the sequence lie sequentially on the same octacycle, starting at $A^{7}$. Lewin (1996) and Cohn (2012), at pp. 156-157, present analyses of this sequence from their own perspectives, but our view is that the characterization of this important progression of minor sixths and dominant sevenths as belonging to a single octacycle on the Tristan tonnetz of Figures 22 and 23 makes for a very satisfactory account of the matter. In fact, the next chord along the octacycle is an $F_{\sharp}^{\,7}$ and we can reasonably take the view that the purpose of Wagner’s intervention of the off-cycle $C_{\sharp}^{\,7}$ just before the would-be $F_{\sharp}^{\,7}$ is to put the brakes on the progression, allowing it to halt at the triadic reduced chord $F_{\sharp}M$ rather than continuing with a fully tetradic $F_{\sharp}^{\,7}$, which would have demanded yet another move to a minor sixth.
The geometrical structure of this progression is particularly striking in the path that it makes as it transverses the configuration in Figure 23. Beginning at the point $A^{7}$ in the lower left corner, the progression progresses in an ever stately way along the line $B_{\flat}m^{6}$, where it lingers, then on to the point $C^{7}$ and then to the line $D_{\flat}m^{6}$ where again it lingers, then on to $E_{\flat}^{\,7}$ and finally to $Em^{6}$, at which point the brakes are applied, and the terminus is reached with $F_{\sharp}M$ substituting for $F_{\sharp}^{\,7}$, as we indicated. Thus we march right around the outer perimeter of the configuration, omitting only the last line – indeed, for had that line been included, then we would be taken back to the beginning. One could even imagine a kind of alternate reality wherein Wagner, in a moment of enthusiasm, decides to repeat the entire sequence to draw out this scene a bit further, dropping the $C_{\sharp}^{\,7}$, playing the full $F_{\sharp}^{\,7}$, then adding in just one more chord – a $Gm^{6}$ – that would take the sequence back to the $A^{7}$ for a repeat, in a kind of perpetuum mobile.  [^31] From this perspective, the progression in this final scene of Götterdämmerung can be understood as a musical realization of a certain quadrilateral in the geometric configuration of the Tristan tonnetz. The music forms a ‘representation’ of the geometry. Perhaps that is the way to think about it.
Interestingly, Wagner uses a version of the same progression based on the same octacycle in Parsifal. Recall the sequence of four minor sixths followed by a diminished seventh at Parsifal’s Act 3 entrance shown in Figure 8, given by
$B_{\flat}m^{6}\to D_{\flat}m^{6}\to G_{\flat}m^{6}\to Em^{6}\to G^{\circ 7}.$ (16)
Now, one can easily check that the ordering of minor sixth chords in general around the perimeter Hamiltonian of the Tristan tonnetz is the same as the ordering of the corresponding triadic minor chords around the perimeter of the Eulerian tonnetz – namely, ascending by fifths counterclockwise. So the conclusions that we reached earlier in our tentative triadic analysis of sequence (16) remain valid in the tetrachordal case, which supports our earlier suggestion, in agreement with Cohn (2012), pp. 142-145, that there is indeed some merit in a reductive approach to Wagner’s tetradic harmonies.
Specifically, we see that the initial minor sixth $B_{\flat}m^{6}$ and the final minor sixth $Em^{6}$ are in polar opposition on the Hamiltonian cycle in Figure 22, as they are in the triadic case, with the first resolution to the ‘Tchaikovsky chord’ $D_{\flat}m^{6}$ halfway between the starting point and the ending point, and the second resolution $G_{\flat}m^{6}$ at a ‘staging post’ one black vertex along the cycle towards the end point. The first, second and fourth chords of (16) lie on the octacycle containing the beginning and ending points. The final resolution to $G^{\circ 7}$, which acts again as a wrap up, is only one semitone away from the would-be continuation $Gm^{6}$, which is another quarter turn around the octacycle.
XI Finale
Back at the Fidelio Cafe, Salviati and Sagredo have paused their conversation to order another bottle. Simplicio rushes in, clutching his coat, shakes the hands of his friends and takes a seat in the empty chair. Sagredo pours him a glass of the wine.
SIMPLICIO – Sorry I am so late! Beastly traffic. Seems like it is always like that in London these days.
SALVIATI – Not to worry, we still have plenty of time.
SIMPLICIO – What is on the program for the concert tonight?
SAGREDO – Wagner, the Wesendonck Lieder. Then the final scene from Götterdämmerung, in the reduction for piano and voice.
SIMPLICIO – Oh, my god! I forgot, I promised Raffaello that I would sing the part of Hagen, the evil half-Nibelung, son of Alberich.
SALVIATI – Good thing you got here eventually. We wouldn’t want to miss that! But Hagen has very little to sing in that last scene. I presume that the soprano will be singing Brünnhilde’s great monologue as she prepares to hurtle herself into the fire?
SIMPLICIO – Yes, and as long as she does not drag me in as well, I should be fine. It is true, Hagen only sings three words, ‘Zurüch vom Ring’. But these words mark the conclusion of the whole of the Ring cycle. And I have to sing a tritone, in the sequence $D_{\flat},B_{\flat},E]$.
SAGREDO – A diminished chord, not easy to get right. And these are the same notes that Gurnemanz, also a bass, sings, in the same key signature, four flats, in Parsifal at the beginning of Act 3, when Parsifal returns as the black knight. Do you think there is some connection between Gurnemanz and Hagen?
SALVIATI – You are always spotting these connections! Well, it’s more likely that Wagner was just recycling some of his ideas. Maybe he was running short of material. But, Simplicio, don’t you have to prepare yourself and get ready in the rehearsal rooms downstairs?
SIMPLICIO – Nonsense, when the time comes, I’ll sing my part just standing here at the table. That is what we agreed with Raffaello. In fact, I may sing it once or twice for you before the concert begins, just to warm up. I look a little like Hagen, wouldn’t you agree?
SALVIATI (gesturing, to change the subject) – While Sagredo and I were waiting for you to arrive, we were discussing the theory of Bolando and Ugopoli.
SIMPLICIO – Oh, that. Well, none of it makes much sense to me. But I love the way they give a tonnetz for the Tristan-genus chords. How did they come up with that?
SALVIATI – Well, that is an interesting story. In an earlier draft of their paper, they introduced their new approach to the tonnetz as a configuration in a purely triadic context. They analyzed passages in Tristan and Parsifal using a reductive approach – that is, using a triadic tonnetz to study tetradic harmonies in the operas. And they made comparisons with works of Chopin and Tchaikovsky, who used some of the same chords. However, not long ago, when they were presenting their work at the University of Southern North Dakota, someone in the audience argued that they needed to justify this reductive approach.
SAGREDO – Not unreasonable, I suppose.
SALVIATI – Also, in the same talk, Bolando and Ugopoli made some remarks to the effect that in their opinion all previous attempts to find a mathematically satisfactory tonnetz theory of tetradic harmonies had been unsuccessful. The audience seemed to be unhappy with this claim, or so it appeared, so Bolando and Ugopoli felt that the only way to resolve the matter – this is what they told me – would be to create a new theory of tetradic harmony based on a tonnetz that was associated with a configuration.
SAGREDO – But who can cook up such theories on demand? Surely, either such a tonnetz exists or it doesn’t and if it does exist, then surely it would have already been found?
SALVIATI – Well, you might have thought so, but the situation was not as simple as that. For sure, a lot of good work has been undertaken by music theorists on the relations between dominant seventh chords and half-diminished sevenths – or minor sixths – but the music theorists would typically take a ‘maximalist’ approach, looking at all possible transitions between these chords (after all, Wagner uses a lot of them), throwing in various other chords as well, for good measure. Just look at Cohn’s book, in the section called, ‘Circumnavigating the Tristan-Genus Universe’. The writing is brilliant, I have to admit, but as a basis for his analysis he uses the Douthett 4-Cube Trio, which is an enormous graph involving twelve dominant seventh chords, twelve minor sixth chords, twelve minor seventh chords, six French sixths,   .  .  .
SIMPLICIO –  .  .  . two turtle doves, and a partridge in a pear tree?
SALVIATI (laughing) – Ha, ha. Yes, well, you see my point.
SAGREDO – You are saying that such a graph, though demonstrably useful, at least in some ways, for descriptive purposes, is unlikely to be recognizable as a mathematical object?
SALVIATI – Precisely.
SIMPLICIO – But what exactly do you mean, in the present context, by ‘recognizable as a mathematical object’?
SALVIATI – Well, that is a difficult question. It is not easy to define what one means by a mathematical object, without going into a great deal of technical detail, but that is not the point. Mathematicians are good at recognizing mathematical objects, even if they can’t explain in simple terms what they mean by the term.
SAGREDO – Are you saying that there are aesthetic elements involved here? Is it a question of ‘mathematical good taste’, whatever that means?
SALVIATI – Well, yes, perhaps, something like that. In part, it is a question of economy. The mathematician will tend to look for the minimal structure that somehow satisfies the conditions they are imposing.
SIMPLICIO – So, that is what Bolando and Ugopoli were doing?
SALVIATI – Well, Simplicio, who knows exactly how they got there. Often, hitting on the right idea is just a matter of good luck or intuition. But from what I understand, they had been looking at the relations between triads when exactly two notes could change in the transition from a major chord to a minor chord. This led to a certain bipartite graph of girth four and degree three that could be unfolded into an Archimedean tessellation of the plane, the famous one of Kepler involving dodecagons, hexagons and squares.
SIMPLICIO – Girth four means that there are tetracycles – these are the squares – and that there is no configuration.
SALVIATI – That is true, but the tessellation can be regarded as a species of tonnetz and it is also a geometric object in its own right. And it gives rise to interesting cycles.
SAGREDO – So, maybe we should not be so dogmatic about what is a tonnetz and what is not. Perhaps we should simply be patient and accept what the mathematics tells us.
SALVIATI – That has always been my view. Follow the flow of the mathematics. Avoid preconceptions. Resist the temptation to prejudge  .  .  .
SIMPLICIO (nodding in agreement) – Yes, of course, preconceptions lead to misconceptions. But you were saying  .  .  .
SALVIATI – Yes, so Bolando and Ugopoli decided to use the same strategy on tetrachords. So, change two notes of the half-diminished seventh in such a way that the result is a dominant seventh. But this leads to a large graph, with twenty-four vertices and six edges meeting each vertex. Unfortunately, there are no configurations that have that structure. In other words – it is not a Levi graph. It admits tetracycles. So what next?
SAGREDO – This is where they added in a further constraint?
SALVIATI – Yes, they added in the constraint that the ‘dissonant’ note in a given tetrachord should be preserved under the transition. This can be done consistently. The flatted seventh in the dominant seventh chord is preserved in the resulting minor sixth, and the sixth of a minor sixth chord is preserved in the resulting dominant seventh. The result, miraculously, is a Levi graph, with no tetracycles, that represents another configuration $\{12_{3}\}$. This time, the D228. And that is how they got their mathematical object.
SIMPLICIO – That’s pretty remarkable. So it seems that the strategy of looking for a configuration paid off.
SAGREDO – So it appears. And they were lucky that those seminar people had somehow spurred them on to think about the problem in the first place.
SIMPLICIO – Yes, they definitely got some good advice. But I guess the real test is to see whether any of the tetradic transitions one sees among the Tristan genus chords in Wagner’s late operas fit the patterns of the cycles of this D228.
SAGREDO – But they do, apparently, that is why it is rather exciting!
SALVIATI – Well, yes, of course, and that is why we came to tonight’s performance, so we could hear with our own ears the octacycle of Bolando and Ugopoli being played as Brünnhilde approaches her immolation.
SIMPLICIO (rising to his feet) – That reminds me, I must rehearse my lines  .  .  .
SALVIATI (gripping the table) – Steady on, mate  .  .  .
SAGREDO (looking the other way) – Perhaps we should change tables  .  .  .
SIMPLICIO (singing loudly) – Zurüch vom Ring! Zurüch vom Ring!
Acknowledgements.
LPH wishes to acknowledge E. Chew, M. Gotham, K. Rietsch and other members of the Music and Acoustics Research Centre (MARC) at King’s College London for stimulating discussions. The authors are grateful for comments made by participants at a MARC seminar in May 2025, and at a colloquium of the Center for Theoretical Physics of the Polish Academy of Sciences in December 2025, where drafts of this work have been presented. We are grateful for the comments and suggestions of two anonymous reviewers, which have led to many improvements. The authors also wish to thank Abdulla Alazemi, J. Armstrong, D. M. Blasius, J. Forth, L. Masiero, R. Morales, P. Nurowski, S. Salamon and others for helpful discussions and correspondence.
APPENDIX: Cycle Count Tables
In this appendix we record the numbers of cycles of various types for the Eulerian tonnetz, the Tristan tonnetz, the Pentatonic tonnetz, the Twelve-tone tonnetz, and the Archimedean tonnetz. The rows are labelled by cycle length and the columns are labelled by the numbers of crossings (the $p$-numbers). Each entry in the table shows the number of cycles of a given length with the given number of crossings. Totals are tabulated for each row and each column. For example, in the case of the Eulerian tonnetz we observe that there are twelve 2pp-hexacycles, four 3pp-hexacycles, and a total of sixteen hexacycles altogether, and that the Eulerian tonnetz admits a grand total of 5409 cycles.
Cycle Count for Eulerian Tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ $11$ Total
66 0 0 $12$ 44 0 0 0 0 0 0 0 0 $16$
88 0 $12$ 0 $24$ 33 0 0 0 0 0 0 0 $39$
$10$ 0 0 $12$ $48$ $60$ 0 0 0 0 0 0 0 $120$
$12$ 0 0 $30$ $72$ $78$ $48$ 0 0 0 0 0 0 $228$
$14$ 0 0 $24$ $24$ $228$ $192$ $120$ $48$ 0 0 0 0 $636$
$16$ 0 0 0 $72$ $168$ $324$ $186$ $156$ 0 0 0 0 $906$
$18$ 0 $12$ $12$ $36$ $108$ $324$ $468$ $348$ $192$ 0 0 0 $1500$
2020 0 0 0 $48$ $42$ $252$ $192$ $312$ $174$ $132$ $30$ 0 $1182$
$22$ 0 0 $12$ 0 $48$ $48$ $84$ $108$ $228$ $144$ $36$ $12$ $720$
$24$ 11 0 0 0 0 0 66 $36$ 33 44 0 $12$ $62$
Total 11 $24$ $102$ $328$ $735$ $1188$ $1056$ $1008$ $597$ $280$ $66$ $24$ $5409$
Table 5: Cycle count for the Eulerian tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for Tristan tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ $11$ Total
66 0 $12$ $12$ 0 0 0 0 0 0 0 0 0 $24$
88 0 0 0 $12$ 33 0 0 0 0 0 0 0 $15$
$10$ 0 0 $12$ $24$ $60$ $24$ 0 0 0 0 0 0 $120$
$12$ 0 0 0 $40$ $78$ $96$ 22 0 0 0 0 0 $216$
$14$ 0 0 0 $72$ $120$ $240$ $192$ 0 0 0 0 0 $624$
$16$ 0 0 $42$ $12$ $144$ $228$ $276$ $144$ 0 0 0 0 $846$
$18$ 0 0 $12$ $72$ $108$ $192$ $364$ $444$ $192$ $40$ 0 0 $1424$
2020 0 $12$ 0 $24$ $54$ $84$ $186$ $372$ $282$ $168$ 0 0 $1182$
$22$ 0 0 $12$ $12$ $12$ $48$ $96$ $120$ $168$ $120$ $24$ $12$ $624$
$24$ 11 0 0 0 0 0 0 0 33 $16$ $18$ $12$ $50$
Total 11 $24$ $90$ $268$ $579$ $912$ $1116$ $1080$ $645$ $344$ $42$ $24$ $5125$
Table 6: Cycle count for the Tristan tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for Pentatonic Tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ Total
66 0 55 $10$ 55 0 0 0 0 0 0 0 2020
88 0 0 $10$ 2020 0 0 0 0 0 0 0 $30$
$10$ 0 55 2020 $45$ $50$ $12$ 0 0 0 0 0 $132$
$12$ 0 55 55 $30$ $60$ $45$ 55 0 0 0 0 $150$
$14$ 0 0 2020 $40$ $120$ $125$ $110$ 55 0 0 0 $420$
$16$ 0 55 0 $25$ $60$ $105$ $80$ $85$ 0 0 0 $360$
$18$ 0 0 55 2020 $30$ $40$ $95$ $90$ $30$ $10$ 0 $320$
2020 11 0 0 0 0 77 55 0 55 55 11 $24$
Total 11 2020 $70$ $185$ $320$ $334$ $295$ $180$ $35$ $15$ 11 $1456$
Table 7: Cycle count for the pentatonic tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for Twelve-Tone Tonnetz
0 11 22 33 44 55 66 77 88 99 $10$ $11$ $12$ $13$ $14$ Total
88 0 55 $30$ $45$ $10$ 0 0 0 0 0 0 0 0 0 0 $90$
$10$ 0 55 55 $30$ $25$ 77 0 0 0 0 0 0 0 0 0 $72$
$12$ 0 0 $10$ $70$ $135$ $80$ 55 0 0 0 0 0 0 0 0 $300$
$14$ 0 55 $40$ $105$ $295$ $365$ $255$ $15$ 0 0 0 0 0 0 0 $1080$
$16$ 0 0 $15$ $85$ $280$ $465$ $530$ $230$ $15$ 0 0 0 0 0 0 $1620$
$18$ 0 55 55 $80$ $350$ $845$ $1085$ $1130$ $320$ 2020 0 0 0 0 0 $3840$
2020 0 0 $35$ $135$ $325$ $1007$ $1685$ $2080$ $1655$ $580$ $22$ 0 0 0 0 $7524$
$22$ 0 55 0 $55$ $245$ $575$ $1415$ $2035$ $2070$ $1410$ $460$ $10$ 0 0 0 $8280$
$24$ 0 55 0 $45$ $155$ $535$ $855$ $1685$ $2345$ $2135$ $1335$ $345$ $10$ 0 0 $9450$
$26$ 0 0 $25$ $25$ $85$ $245$ $615$ $895$ $1390$ $1610$ $1410$ $965$ $285$ $10$ 0 $7560$
$28$ 0 0 0 $25$ 55 2020 $105$ $140$ $150$ $255$ $305$ $200$ $170$ $60$ 55 $1440$
$30$ 11 0 0 0 0 77 0 $15$ 2020 2020 $16$ $10$ $30$ 2020 55 $144$
Total 11 $30$ $165$ $700$ $1910$ $4151$ $6550$ $8225$ $7965$ $6030$ $3548$ $1530$ $495$ $90$ $10$ $41400$
Table 8: Cycle count for twelve-tone Tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.
Cycle Count for the Archimedean Tonnetz
0 11 22 33 44 55 Total
44 0 0 33 0 0 0 33
66 0 66 $12$ 22 0 0 2020
88 0 66 0 $18$ 0 0 $24$
$10$ 0 0 $12$ $18$ $12$ 66 $48$
$12$ 11 0 0 22 33 66 $12$
Total 11 $12$ $27$ $40$ $15$ $12$ $107$
Table 9: Cycle count for the Archimedean tonnetz. The rows are labelled by cycle length and the columns are labelled by the number of $p$-crossings.